### v1

# Human or LLM as Standardized Patients? A Comparative Study for Medical Education

Abstract
Standardized Patients (SP) are indispensable for clinical skills training but remain expensive, inflexible, and difficult to scale. Existing large-language-model (LLM)–based SP simulators promise lower cost yet show inconsistent behavior and lack rigorous comparison with human SP. We present EasyMED, a multi-agent framework combining a Patient Agent for realistic dialogue, an Auxiliary Agent for factual consistency, and an Evaluation Agent that delivers actionable feedback. To support systematic assessment, we introduce SPBench, a benchmark of real SP–doctor interactions spanning 14 specialties and eight expert-defined evaluation criteria. Experiments demonstrate that EasyMED matches human SP learning outcomes while producing greater skill gains for lower-baseline students and offering improved flexibility, psychological safety, and cost efficiency.
Human or LLM as Standardized Patients?
A Comparative Study for Medical Education
Bingquan Zhang1,2, Xiaoxiao Liu2, Yuchi Wang2, Lei Zhou3, Qianqian Xie1†, Benyou Wang2† 1 School of Artifical Intelligence, Wuhan University 2 The Chinese University of Hong Kong, Shenzhen 3 Freedom AI
1 Introduction
Medical education underpins professional healthcare training, fostering essential competencies in clinical reasoning, communication, and decision-making cleland2022researching. Deficiencies in these skills contribute to diagnostic errors and preventable harm, with communication failures alone responsible for thousands of in-hospital deaths annually ball2015improving. To improve learning efficiency and scalability, recent studies have explored artificial intelligence (AI)—particularly LLM—for clinical education. LLM enable interactive simulation and automated feedback at low cost, offering new possibilities for scalable, practice-oriented training  thirunavukarasu2023large.
Patient simulation is central to such experiential learning, allowing students to practice reasoning and communication in realistic but safe settings. The Standardized Patient—actors trained to portray clinical cases—is widely recognized as the gold standard  sayers2024database; ma2023standardized, yet remains costly, labor-intensive, and difficult to scale, limiting case diversity and accessibility zendejas2013cost.
To mitigate these constraints, previous researchers have explored LLM-based virtual SP simulators as low-cost and scalable alternatives ye2025multimodal; du2024llms; khandwaha2025designing. Most systems adopt a single-agent, prompt-driven architecture, where a pretrained model performs patient simulation, intent understanding, and dialogue control within one process. This unified design couples persona control, information disclosure, and evaluation, leading to unstable role behavior, factual errors, and uncontrolled information flow sirdeshmukh2025multichallenge; laban2025llms. Evaluation methods are usually ad hoc, based on coarse rubrics or single-turn scoring, which limits analysis of interactive reasoning and learning outcomes burke2024assessing; wang2025survey. As a result, whether virtual SP can match the performance and educational value of human SP remains unclear simzine2025vsp.
In this research, we introduce EasyMED, a multi-agent framework that transforms virtual SP training into a structured and interactive process. Unlike traditional single-agent models li2023chatdoctor, EasyMED decomposes patient simulation, intent recognition, and evaluation into coordinated agents with distinct roles. Each agent aligns with the workflow of real clinical practice: the Patient Agent simulates realistic symptom expression and emotional behavior, the Auxiliary Agent interprets the doctor’s conversational intent to guide contextually appropriate responses, and the Evaluation Agent monitors and assesses interactions to ensure consistency and clinical validity. This design establishes an interpretable framework that enhances controllability and error traceability through clearly defined, task-specific agent roles.
To benchmark interactive behaviors that matter pedagogically, we introduce SPBench, a dataset derived from authentic SP-student dialogues covering 14 specialties and eight expert-defined assessment criteria. It standardizes prompts, information disclosure rules, and scoring procedures, enabling consistent comparison of models and prompting strategies under realistic clinical scenarios.
We further validate EasyMED through a four-week controlled study comparing it with human SP training under identical content and scoring rubrics. The results demonstrate that EasyMED achieves comparable or even superior learning outcomes—particularly among novice learners—while providing greater flexibility, psychological safety, and cost efficiency.
In summary, our contributions are threefold: (1) Multi-Agent Framework for Virtual SP Training: We propose EasyMED, effectively simulates the structured interactive process of the realistic SP and training doctors with three coordinated agents. This structure enables controllable, persona-stable interactions and provides transparent, multi-dimensional feedback. (2) Benchmarking with SPBench: We construct a benchmark from authentic SP–student dialogues covering 14 specialties and eight evaluation criteria. It standardizes prompts, disclosure rules, and scoring, supporting reproducible, turn- and session-level evaluation of interactive behaviors. (3) Empirical Validation of Educational Impact: We conduct a four-week controlled study comparing EasyMED with human SP training under identical test content and rubrics, demonstrating comparable learning outcomes, improved accessibility, and clear advantages for lower-baseline learners.
2 Background and Motivation
2.1 The Definition of Standardized Patients
Definition
Standardized Patients (SP) is a training approach in which individuals are trained to consistently simulate the history, personality, physical findings, and emotional states of real patients. SP is widely used in medical education, especially in Objective Structured Clinical Examinations (OSCE), providing a safe and repeatable environment for practicing key clinical skills such as history taking, physical examination, and communication cook2009virtual; berman2016role; kononowicz2015virtual. This allows students to develop these abilities without risking harm to real patients.
Challenge
While SP is effective, it faces several challenges, including high costs for recruiting and training actors, logistical constraints from fixed locations and schedules, and inconsistent performance due to factors like fatigue or mood hamilton2024evolution; kuehne2018standardized. These limitations restrict the scalability and accessibility of SP-based training, particularly in resource-limited settings.
2.2 LLM Simulator vs. Standardized Patient
As discussed in Section 2.1, the high cost, limited scalability, and inconsistent performance of traditional SP have prompted researchers to explore LLM-based virtual SP as an alternative. While LLM-based SP reduce costs, improve scalability, and standardize responses compared with human SP, they still show unstable behavior, lack structured feedback, and have limited validation against human SP. Building on these questions, our EasyMED framework further enhances virtual SP capabilities by introducing a multi-agent design, structured case control, and integrated evaluation modules.
Metric Human SP Existing LLM-based SP Our EasyMED
Cost High actor, training and venue costs Low API costs; limited assessment Low costs with built-in multi-dimensional assessment
Time Fixed scheduling slots Instant but single-agent Instant with multi-agent coordination
System Architecture Single-agent role-play; no module separation wang2023notechat; wang2024patient Multi-agent with Patient, Assistant, Evaluation
Assessment Capability Manual checklists Manual or simple scoring; no integrated module li2024leveraging Built-in Evaluation Agent with automated, multi-dimensional feedback
Benchmarking No real-case benchmark; mostly synthetic sardesai2024utilizing; ye2025multimodal SPBench: real SP–student dialogues (12 specialties, 8 dimensions)
Empirical Validation Local experience only Few or no long-term comparative studies ye2025multimodal Four-week study comparing EasyMED vs. Human SP
Persona Stability Actor variability Persona inconsistency, leakage, harmful replies du2024llms Prompt control + multi-agent improves stability
Consistency Performance varies across actors Standardized but unstable persona Multi-agent design improves persona stability
Feedback Manual scoring Limited or no automated feedback Evaluation Agent + SPBench for reliable feedback
Reproducibility Limited, hard to repeat Anytime but limited diversity Anytime with curated real-case benchmark
Psychological Safety Higher performance anxiety Lower pressure vs. Human SP almansoori2025self. Lower pressure + structured, transparent assessment
Table 1: Integrated comparison of Human SP, existing LLM-based SP systems, and our EasyMED framework across operational efficiency, system design, teaching quality, and learner experience.
Figure 1: The multi-agent architecture of our virtual medical training system. This “division of labor” model includes: (1) a Patient Agent to simulate patient dialogue, (2) an Intention Identification Agent to interpret the user’s clinical queries, and (3) an Evaluation Agent to assess performance against ground truth.
As shown in Table 1, these limitations are evident when comparing Human SP, existing LLM-based SP systems, and our EasyMED framework across operational efficiency, system design and validation, teaching quality, and learner experience.
3 Methodology
3.1 EasyMED: Multi-Agent Framework
Building on the limitations of single-agent simulators discussed in Section 1, EasyMED is designed as a controllable and modular training framework. Existing virtual SP systems integrate patient simulation, intent understanding, and evaluation within one model, which often leads to persona drift, uncontrolled information disclosure, and weak or non-actionable feedback du2024llms. To address these issues, EasyMED adopts a multi-agent division-of-labor architecture (Figure 1) that explicitly separates patient simulation, intent interpretation, and evaluation, while maintaining coordination within a single session.
During a training session, a student’s utterance is first analyzed by an Auxiliary Agent, which identifies the clinical intent behind the question and maps it to one of 31 expert-defined intent categories (see Appendix 9 for qualifications). This output conditions the Patient Agent, which generates clinically grounded responses from curated case scripts under explicit information-disclosure constraints to preserve realism. All interactions are then logged and evaluated by an Evaluation Agent, which compares the student’s inquiries with predefined checklists and applies an LLM-based judge to assess completeness and reasoning accuracy. At the end of each consultation, the Evaluation Agent produces structured feedback summarizing strengths and areas for improvement.
3.2 SPBench: Benchmark for Virtual SP
To enable systematic and reproducible evaluation of LLM-based SP, we constructed SPBench, a benchmark dataset derived from authentic SP–physician dialogues.
Figure 2: Distribution of clinical cases in the SPBench dataset by medical department.
Case Source and Scope
We collected 58 real clinical cases. From the doctor-patient dialogue records associated with these cases, we processed and extracted a dataset containing 3,208 question-and-answer pairs. The source materials include the Manual for Writing Standardized Patient Cases [^1] https://weread.qq.com/web/bookDetail/1a432e10811e38145g0194ab and the A Practical Tutorial for Standardized Patients [^2] https://www.pumpedu.com/home-shop/7125.html. All records were fully anodized and used solely for academic research. They were also reviewed by experienced clinicians (see Appendix 9 for expert qualifications) to ensure accuracy and representativeness. As shown in Figure 2 and Figure 3, the specialty distribution of SPBench spans 14 clinical departments. General Surgery is the most represented specialty, accounting for 13.79% of the cases, followed by other departments such as Cardiology, Respiratory Medicine, and Psychiatry. The age of the patients ranges from children to the elderly, with the majority concentrated in the 21-60 age group. Regarding gender, the dataset is composed of 35 female cases, representing 60.3% of the total, and 23 male cases, representing the remaining 39.7%.
Figure 3: Demographic distribution of cases in the SPBench dataset. The left panel shows the age distribution, and the right panel shows the gender distribution.
Data Structure
SPBench provides two core types of content: 1) Case Scripts: These provide standardized patient histories and physical examination findings for each case; and 2) Authentic Dialogue Records: These are turn-by-turn transcripts of conversations between physicians and human SP. The records serve as a benchmark to compare the conversational quality of an LLM-based virtual SP against a human SP for the same case.
Quality Control
To ensure accuracy and reliability, all dialogue transcripts were digitized from the two books in Section 3.2 using Optical Character Recognition (OCR) and then manually proofread by three fifth-year medical students. They corrected OCR-induced errors such as typos and punctuation to ensure fidelity to the original source material. This process preserves the authenticity and consistency required for benchmarking.
3.3 Evaluation Protocol
To systematically assess the performance of LLM as Standardized Patients, we (i) define what to measure via eight core evaluation criteria; (ii) validate the reliability of our automated scoring against clinical experts; (iii) conduct fair comparisons between human SP, EasyMED, and an external framework under identical inputs and scoring; and (iv) perform component-level ablations within EasyMED to analyze where gains come from. The following subsections detail each component.
Evaluation Metric
In collaboration with three clinical experts (see Appendix 9 for qualifications), we defined eight core dimensions for evaluating the virtual patient’s response quality, as detailed in Table 3. During the evaluation, experts rated performance for each dimension on a 5-point scale. These scores were then linearly transformed to a 100-point scale for reporting to enhance the granularity and readability of the results.
Expert Rating and Correlation Analysis
Considering the high cost of large-scale manual annotation, we first used a well-designed evaluation prompt (Appendix I.4) and employed GPT-4o as an automated evaluator to score all results. To verify its reliability (see Appendix 9 for qualifications), we randomly selected 30 samples for blind review by two clinical experts who were unaware of the GPT-4o ratings. The automated scores correlated strongly and significantly with the experts’ average ratings (Pearson’s $r=0.79$; Table 2). Although GPT-4o mean score (78.5) was consistently stricter than the experts’ mean (83.9), this high alignment supports its use as a reliable proxy for large-scale evaluation.
Automated Professional A Professional B
Average Score 78.5 86.3 81.5
Standard Deviation 8.8 9.3 9.8
Correlation 0.79
Table 2: Inter-Rater Reliability between an Automated Evaluator and Human Experts
Evaluation Criterion Description
Query Comprehension Assesses the ability to accurately understand the physician’s questions, grasp the core intent, and provide relevant answers without misinterpretation.
Case Consistency Evaluates whether responses are factually consistent with the predefined patient case, including symptoms, medical history, and key facts. Checks for contradictions or information not present in the case.
Controlled Disclosure Measures the ability to provide information passively and appropriately, answering only what is asked without volunteering unsolicited details. Focuses on conciseness and relevance.
Response Completeness Assesses whether responses are comprehensive and address all parts of a physician’s query, avoiding the omission of crucial information from the case profile.
Logical Coherence Evaluates the internal logical consistency of the information provided. Ensures that reported symptoms and their characteristics (e.g., location, duration) are coherent from a patient’s perspective.
Language Naturalness Assesses whether the communication style is natural and patient-like, using plain language and avoiding overly technical medical jargon. Focuses on the authenticity of the linguistic style.
Conversational Consistency Measures the ability to maintain consistent information across a multi-turn dialogue. Checks for self-contradictions and ensures stability throughout the interaction.
Patient Demeanor Evaluates the ability to maintain a patient-like demeanor, demonstrating appropriate patience, cooperation, and emotional tone, even when faced with repetitive or clarifying questions.
Table 3: Definitions of the eight evaluation criteria for assessing the quality of the conversational agent.
Comparison between human and LLM-based SP
To compare the responses of LLM-based SP with those of a human SP under identical context, we first isolated the physician’s questions from the dialogue records in SPBench. We then used these exact questions and patient case as inputs for the LLM-based SP. This method allowed us to directly compare the responses generated by our virtual patients against those of the human SP under identical conditions.
Comparison with External Frameworks
Beyond comparing against human SP under identical prompts, we further benchmarked our system against an open-source LLM-based SP framework, EvoPatient du2024llms. Using the same SPBench physician questions and case profiles, and the same automated evaluator and scoring protocol, basic settings are summarized in Appendix F. we report the system-level scores in Table 4.
System Configuration Score
Human SP 97.33
EasyMED (ours) More Agent 96.98
EvoPatient Authors setup 93.33
Table 4: Comparison of different frameworks on SPBench under the same evaluation protocol.
Ablation on EasyMED Components
We conducted an ablation study to quantify the contribution of key components in our virtual SP design. For each model, we started from a baseline configuration and incrementally added components: (1) Baseline no chain-of-thought or intent recognition; (2) CoT adds explicit step-by-step reasoning to the baseline; (3) Intent Recognition further adds an intent-recognition module on top of CoT.
Model Name Baseline COT Intent Recognition
Qwen3-8B 81.12 81.95 81.64
Qwen3-32B 88.87 91.89 94.08
DeepSeek-R1 93.11 94.90 95.72
GPT-4o 89.54 90.08 93.70
Gemini 2.5 Pro 95.04 95.77 96.98
Table 5: Effect of CoT and intent recognition modules across different models.
Table 4 shows that EasyMED matches human SP performance (96.98 vs. 97.33) and clearly outperforms the external framework EvoPatient (93.33). These results indicate that our multi-agent design and evaluation-aligned prompting produce higher-fidelity SP behaviors under identical inputs and scoring.
As shown in Table 5, adding COT to the baseline improves most models, and layering intent recognition on top yields further gains for medium-to-large models (e.g., Qwen3-32B, DeepSeek-R1, GPT-4o, Gemini 2.5 Pro). However, these advanced components are capacity-sensitive: for a small model such as Qwen3-8B, intent-guided CoT can introduce over-constraints and slightly reduce the score relative to plain CoT.
Group Sequence Mean Score (±SD) at Test Point Mean Score Gain (±SD) by Phase
Pre-test Mid-test Post-test Phase 1 (Wks 1-2) Phase 2 (Wks 3-4) Total Gain
[…]
(N=7) (±11.45) (±6.88) (±4.53) (±7.82) (±3.65) (±7.45)
[…]
(N=7) (±13.04) (±7.42) (±4.93) (±7.45) (±3.25) (±8.36)
Table 6: The table presents mean scores at each test point and the corresponding mean score gains during each training phase. Participants were in either Group A or Group B. All values are mean $\pm$ standard deviation.
4 Experiment Protocol
After validating our virtual SP on SPBench, we conducted a user study to evaluate its educational impact in a real-world setting. This section describes the four-week randomized crossover trial with undergraduate medical students, designed to test whether EasyMED outperforms traditional human SP training.
Study Period Timeline Group A Group B
Baseline Week 0 Pre-Test Pre-Test
Period 1 Weeks 1-2 Intervention A: Intervention B:
EasyMED Training Human SP Training
End of Week 2 Mid-Test Mid-Test
Period 2 Weeks 3-4 Intervention B: Intervention A:
Human SP Training EasyMED Training
End of Week 4 Final Test & Questionnaire
Table 7: Design of the Experiment, illustrating the sequence of interventions (EasyMED Training and Human SP Training) and assessments for Group A and Group B.
4.1 Experimental Design
We ran a four-week randomized crossover experiment comparing EasyMED with traditional human SP training. Each participant experienced both training methods in different phases, enabling within-subject comparisons and controlling for individual differences in baseline competency. The procedure is shown in Table 7.
Patient Cases
During the intervention, students practiced with 20 standardized clinical cases covering core diseases at the scope and difficulty of the National Medical Licensing Examination (see Appendix D).
4.2 Participants
We recruited 20 fourth- and fifth-year undergraduate medical students, all of whom completed a pre-test to assess baseline knowledge and received a 10-minute EasyMED orientation. Students with scheduling conflicts or extreme pre-test scores were excluded, resulting in a final sample of 14 (7 males, 7 females; mean age 23, range 21–-24). All had completed core clinical coursework but had not yet taken the National Medical Licensing Examination.
To ensure baseline equivalence, participants were rank-ordered by pre-test scores and alternately assigned to two sequence groups. Three experienced SP professionals delivered the human SP sessions. All participants received hourly compensation in accordance with ethical regulations.
5 Results and Analysis
This section presents and analyzes the results from Section 4, focusing on whether the Virtual SP-based EasyMED can deliver learning outcomes equivalent to or better than traditional human SP training.
5.1 Evaluating Overall Improvement via OSCE
Figure 4: Boxplots comparing the baseline scores of Group A and Group B. Each circle represents the score of an individual participant. The similarity in the distributions demonstrates that the two groups were comparable at the start of the experiment.
Figure 5: Learning trajectories of individual students and group averages for Group A (left) and Group B (right) across three assessment points. Each colored line tracks an individual student’s performance, while the bold dashed line represents the group’s average score.
Both groups started from an equivalent baseline We first confirmed that the two groups were comparable at baseline. Baseline equivalence. Figure 4 shows that the two groups had similar initial score distributions. An independent-samples t-test confirmed no significant difference between Group A (mean = 70.56) and Group B (mean = 69.84), t(12) = 0.16, p = 0.88, indicating comparable baselines.
Overall Performance Improvement:
Both training methods produced clear, comparable gains. Specifically, As shown in Table 6, Group A’s average score rose from 70.56 to 87.44 (+16.89), while Group B’s increased from 69.84 to 85.20 (+15.36). Figure 5 illustrates this upward trend across sessions.
Figure 6: Comparison of mean score gains in Phase 1 for the Human SP and EasyMED training methods. The participants are stratified into low- and high-performing groups based on their pre-test scores. Error bars indicate the standard error of the mean.
Improvement at Each Stage
Most gains occurred in the first two weeks, followed by a slowdown in Weeks 3-–4. Table 6 shows that Group A improved by 15.51 points and Group B by 12.17 points in the first two weeks. In Weeks 3–4 the rate of improvement slowed markedly, indicating diminishing returns and underscoring the strong initial impact of both training methods.
Figure 7: This chart shows the total score gain for each group, broken down by training phase. The total height of each bar represents the average total gain for that group. The stacked segments show the portion of the gain that occurred during Phase 1 (Weeks 1-2) and Phase 2 (Weeks 3-4).
Figure 8: Comparative heatmap of the weekly practice time distribution for the EasyMED and Human SP groups. The left panel shows the EasyMED group, and the right panel shows the Human SP group. In both heatmaps, the x-axis represents the day of the week, and the y-axis represents the time of day. The color intensity and the white number in each cell indicate the number of students who practiced during that time slot.
Improvement by Skill Level
EasyMED delivered the largest benefits for low-baseline students early on. We split participants into high-baseline (top three) and low-baseline (bottom four) subgroups based on pre-test scores. As shown in Figure 6, low-baseline students using EasyMED gained an average of 21.83 points versus 16.58 in the human SP group. High-baseline students improved more modestly—7.10 for EasyMED versus 6.30 for human SP—showing that EasyMED is particularly advantageous for learners starting at a lower skill level.
Improvement by Training Method
Across the crossover design, EasyMED produced larger score gains than traditional SP training. Although the final post-test scores for Group A (M = 87.44) and Group B (M = 85.20) were not significantly different ($p>.05$), the crossover design allowed us to analyze the score gains from each training method independently (Figure 7). Training with EasyMED led to an average gain of 9.35 points, compared with 6.77 points for training with a human SP. This indicates that, while both methods are effective, EasyMED yields greater skill improvement than traditional SP training.
5.2 Behavioral Analysis via Survey and Logs
To investigate why EasyMED is effective, we analyzed both behavioral logs and questionnaire responses.
Training Flexibility
Students rated EasyMED more usable and used it far more flexibly, with practice extending into evenings and weekends. Students gave EasyMED significantly higher usability scores than human SP (4.8 vs. 3.5, $p<.01$). This preference was reflected in activity patterns (Figure 8): human SP sessions clustered during weekday working hours, whereas EasyMED practice was spread throughout the day and peaked on evenings and weekends. These patterns indicate that EasyMED overcomes the scheduling limits of traditional training, offering more flexible and self-paced learning.
Psychological Safety and Engagement
EasyMED provided a lower-anxiety environment that encouraged deeper engagement and more questioning. Students reported significantly less learning anxiety with EasyMED than with human SP (mean score 2.5 vs. 3.2, $p<.01$). Interviews echoed this, with participants noting they were “not afraid to make mistakes” and felt comfortable “asking follow-up questions repeatedly.” This sense of safety translated into higher engagement: sessions with EasyMED averaged more dialogue turns (54 vs. 47) and longer durations (28:49 vs. 15:17) than human SP sessions (Table 8). Although part of the extra time reflects typing, the higher number of dialogue turns more strongly indicates engagement, suggesting that the low-anxiety environment promotes deeper reasoning and inquiry.
Metric EasyMED Human SP
Student Engagement
Learning Anxiety Score [^1] Anxiety was rated on a scale where lower scores are better. The difference is statistically significant ($p<.01$). 2.5 3.2
Average Dialogue Turns 54 47
Average Interaction Duration 28m 49s 15m 17s
Cost-Effectiveness
Per-Session Cost \$0.725 \$52.95
Table 8: Comparison of student engagement and cost-effectiveness metrics between EasyMED and human SP.
Cost-Effectiveness:
EasyMED delivers comparable training at a fraction of the per-session cost of human SP, making large-scale deployment feasible.
Beyond its pedagogical benefits, EasyMED offers a substantial advantage in scalability and cost. As shown in Table 8, the per-session cost for an EasyMED interaction is approximately \$0.73, compared with \$52.95 for a human SP. This dramatic reduction in operational expense highlights the potential for EasyMED to provide high-quality clinical training at scale at only a fraction of the cost of traditional methods.
6 Conclusion
This study examines the use of large language models in SP training. We present EasyMED, a multi-agent framework that separates patient simulation, intent recognition, and evaluation into coordinated agents aligned with real clinical workflows. To enable consistent evaluation, we build the SPBench benchmark from SP–student dialogues across 14 specialties and expert-defined criteria. A four-week controlled study shows that EasyMED matches human SP training in learning outcomes and improves performance for novice learners through higher flexibility, safety, and lower cost.
Our work demonstrates the feasibility of multi-agent, LLM-based SP systems for scalable and effective clinical training. Future work will test larger cohorts, integrate multimodal interaction, and study long-term learning effects.
7 Limitations
Our study has several limitations. It was conducted at a single institution with a relatively small and homogeneous cohort, so broader validation across different settings and learner populations is needed. In addition, EasyMED currently supports only text-based interactions without non-verbal or multimodal cues, which are important for authentic clinical communication. Finally, although our automated scoring showed strong correlation with expert ratings, it may still overlook subtle aspects of dialogue quality and learner behavior. Future work will include larger multi-site and longitudinal studies, integration of multimodal interaction channels, and refinement of evaluation metrics to better capture nuanced performance.
8 Ethical Statement
The study was approved by the institute on August 30, 2025. All annotators were fairly compensated, adhering to the standard hourly wage practices of their respective states.
Appendix A Related Work
This section reviews research on virtual patients and intelligent tutoring, the use of large language models (LLM) in education, and automated assessment of complex, interactive skills. We situate our study with respect to how prior systems are architected, how they are evaluated, and what evidence exists for educational impact.
A.1 Virtual Patients and Intelligent Tutoring
Virtual patients (VP) have long supported safe practice of diagnostic reasoning and communication in medical education cook2009virtual; berman2016role; kononowicz2015virtual. Early systems were primarily rule- or script-based, providing structured but inflexible interactions and limited behavioral realism huang2007virtual. Recent work explores LLM-driven VP to increase linguistic fluency and adaptability du2024llms; almansoori2025self; grevisse2024raspatient; steenstra2025scaffolding. However, most LLM-based VP adopt a paradigm that collapses patient simulation, dialogue control, and assessment into one model, typically focusing on the consultation/history-taking phase while leaving evaluation ad hoc and feedback coarse. Architecturally, this design makes persona stability and information disclosure emergent rather than governed; methodologically, it leaves limited support for actionable, multi-dimensional feedback. Our work differs by factorizing the interaction into coordinated agents with first-class interfaces for learner intent, disclosure policy, and inquiry coverage, and by pairing the simulator with a protocol that evaluates turn- and session-level behaviors relevant to clinical training.
A.2 Large Language Models in Education
LLMs have been studied for tutoring, content generation, and role-playing across domains gan2023large; wang2024large; xu2024large; safranek2023role. In medical education, they have been used to generate patient histories and to support clinical decision making, and to simulate patient dialogues for practicing history-taking waisberg2024large; thirunavukarasu2023large; almansoori2025self; fan2023nphardeval. Persistent challenges include factual reliability, long-context consistency, and alignment with professional standards and safety constraints li2024enhancing; mahaut2024factual; gai2024achieving. Most studies report surface metrics or static outcomes (e.g., diagnosis/referral accuracy) rather than interaction competencies. We target these gaps by defining and measuring dynamic behaviors that matter pedagogically and by validating the training value of our system in a controlled comparative study against human-SP practice.
A.3 Automated Assessment of Complex Skills
Automated assessment with LLMs has advanced in essay scoring and feedback zhangpersonaeval; lee2024analyzing; gusev2024pingpong and program analysis for coding education parvathy2025automated; chen2020analysis, but much of this work provides single-dimensional scores and limited interoperability with respect to process. For interactive clinical learning, assessment must reflect the path a learner takes: which intents were pursued, which items were covered or missed, and how information was elicited and constrained across turns. Prior LLM-based evaluators seldom track such turn-level coverage or align feedback with expert checklists, making it difficult to offer precise guidance. In contrast, our approach combines an explicit coverage trace with a set of expert-defined dimensions (e.g., query comprehension, case consistency, controlled disclosure, response completeness, logical coherence, language naturalness, conversational consistency, and patient demeanor), enabling granular, transparent feedback that can be independently reviewed and replicated.
Appendix B Data Annotation Statement
The annotation and validation of all datasets in this study were conducted by qualified medical professionals. Three clinical experts reviewed all cases to ensure medical accuracy and representativeness. Two licensed physicians with outpatient referral experience independently annotated and verified the dialogue data according to predefined guidelines, with only consensus cases retained. In addition, three fifth-year medical students performed manual proofreading and transcription of dialogue records to correct OCR errors. For the creation of standardized patient scripts, three SP education specialists collaborated with three professional standardized patient actors to design and validate clinically authentic dialogues. All contributors participated voluntarily and provided informed consent under institutional ethical approval.
Appendix C Construction of the Intent Recognition Test Dataset
To ensure a rigorous and accurate evaluation of the models’ intent recognition capabilities, we constructed a high-quality test dataset. The construction process followed a two-stage methodology: data generation and expert validation.
Data Generation
We began with a predefined framework of 31 core clinical intents. Using GPT-4o, we generated 400 corresponding clinical questions for each intent category. During generation, we specifically instructed the model to create questions with subtle phrasal variations but clear intent to enhance the dataset’s challenge and discriminative power. This stage yielded a preliminary corpus of 12,400 questions.
Expert Validation and Curation
The preliminary corpus was subsequently reviewed by a three-person expert panel, composed of professional medical personnel. The panel’s task was to remove any questions that were ambiguous, clinically unrealistic, or had unclear intent attribution to ensure the high quality and validity of each entry in the final dataset. After meticulous manual filtering and proofreading, we finalized a validated dataset containing 4,631 clinical questions.
Result
As shown in Table 10, we evaluated several mainstream models on our constructed dataset. The results clearly indicate that Gemini2.5-flash performed best among all models, achieving an accuracy of 96.3% and a macro-average F1-score of 95.0%, significantly outperforming other baseline models. Based on this superior performance, we selected Gemini2.5-flash as the core intent recognition model for the EasyMED system to ensure accurate interpretation of learner input in complex clinical interactions.
Table 9: The structured framework of inquiry intents for clinical history taking. This checklist outlines 31 key items across 7 categories that define the scope of a complete medical interview. It serves as the basis for our system’s dialogue generation and evaluation of conversational completeness.
No. Category Question Items
Patient Identification
1 Demographics Name, Age, Gender, Occupation
Chief Complaint & Present Illness
2 Symptoms Chief complaint
3 Onset Time of symptom onset
4 Cause Precipitating factors
5 Location Site of the symptom
6 Character Characteristics of the symptom
7 Duration Duration and frequency
8 Modifiers Exacerbating/relieving factors
9 Associated Associated symptoms
10 Progression Disease progression
11 Treatment Previous treatments and outcomes
12 Tests Previous investigations and results
System Review
13 General Mental status, sleep, and appetite
14 Elimination Urinary and bowel habits
15 Changes Weight changes and energy levels
Past Medical History
16 Health General health history
17 Chronic Hypertension, Diabetes, CAD
18 Infectious Hepatitis, Tuberculosis
19 Surgical Operations and trauma
20 Transfusions Blood transfusion history
21 Allergies Drug and food allergies
22 Immunization Vaccination history
Personal & Social History
23 Travel Residence and travel history
24 Habits Tobacco, alcohol, substance use
25 Occupation Work environment and exposures
26 Sexual High-risk sexual behaviors
Family & Gynecological
27 Obstetric Marital and obstetric history
28 Family Family medical history
29 Menstrual Menstrual history (female)
Additional Items
30 Communication Small talk and patient education
31 Other Other relevant inquiries
Note: CAD = Coronary Artery Disease
Model Name Accuracy (%) Macro Average (%)
Precision Recall F1-Score
ChatGLM4.5 89.6 95.2 89.5 92.3
Qwen3-8B 86.5 86.4 86.5 86.4
Qwen3-32B 89.6 93.6 89.6 91.5
GPT-4o 92.6 95.9 92.6 94.2
DeepSeek-V3 87.1 92.2 87.1 89.5
Gemini2.5-flash 96.3 96.1 93.9 95.0
Table 10: Performance comparison of different models on the balanced intent recognition task, reflecting realistic fluctuations. All scores are percentages (%).
Appendix D Clinical Case Data Preparation
This section details the source and selection criteria for the 20 clinical cases used in our user study, as well as the data preparation process for both the human SP and the EasyMED system.
D.1 Case Source and Selection
A panel of medical experts selected 20 clinical cases from the authoritative "Peking Union Medical College Hospital Clinical Thinking Training Case Collection," ensuring they were aligned with the curriculum for fifth-year undergraduate medical students. The distribution of these cases across demographics and medical specialties is illustrated in Figure9.
Figure 9: Distribution of cases
The dataset is evenly balanced with 10 male and 10 female patients. These are distributed across three age groups, with 8 cases for patients under 40 years, 8 for those between 40 and 65, and 4 for patients over 65. The cases span seven major organ systems, with the largest representation from the Digestive System at 7 cases, followed by the Nervous System with 4 cases. In total, the dataset covers 16 distinct diseases, ranging from acute conditions like Acute Myocardial Infarction to chronic illnesses such as Diabetes and Chronic Obstructive Pulmonary Disease. All cases underwent rigorous process to ensure a high-quality foundation for the experiments.
D.2 Data Processing Workflow
To support the EasyMED multi-agent framework in achieving high-fidelity SP simulations, we designed a two-stage data processing workflow: 1) generating performable dialogue scripts for human SP; and 2) structuring the clinical cases to be compatible with LLM inputs, aiming to reduce the risks of hallucination and irrelevant responses, and to support the full-workflow clinical simulation and automated evaluation.
Human SP Script Generation
In the process of creating high-fidelity SP scripts, we invited three SP education experts to collaborate with three professional SP actors. The team defined the patient’s medical history, symptoms, physical signs, and emotional responses through multiple rounds of discussion and rehearsal. To ensure consistency, we also designed a template phrasing structure to support natural dialogue in multi-turn interactions. Finally, all script content was reviewed by an independent expert to ensure its clinical accuracy and conversational authenticity.
LLM Input Case Structuring
To enable the LLM to accurately understand and adhere to the case settings, the original text-based cases needed to be converted into a structured format. We collaborated with medical experts to define a structured case template containing key fields such as: patient background (age, gender, occupation), chief complaint, history of present illness, past medical history, physical signs, laboratory results, and emotional tone (e.g., anxious, calm). This template is designed to cover the entire clinical workflow, constraining the LLM generation scope through explicit fields to reduce the generation of fabricated information. We utilized the GPT-4o model, combined with custom prompt engineering, to automatically map the unstructured case text to the predefined fields. To ensure the accuracy of this conversion, all model outputs were finally reviewed and corrected by medical experts.
Appendix E Outcome Measures
We collected data through both quantitative and qualitative methods. Our primary and secondary outcome measures are detailed below.
E.1 Primary Outcome Measure: OSCE Scores
We administered OSCE tests to all students at three time points: pre-experiment, mid-experiment, and post-experiment. To avoid learning effects, the cases used in the three tests were different but were reviewed by experts to ensure consistent difficulty and assessment points. Scoring was performed independently by two blinded examiners who were unaware of the students’ group assignments, ensuring objectivity.
E.2 Secondary Outcome Measure: Subjective Questionnaire
At the end of the experiment, we used a subjective questionnaire with 25 items across four dimensions (Usability, Authenticity, Learning Value, and Learning Anxiety) to collect students’ perceptions and experiences of the two training modalities.
E.2.1 Part 1: Background Information
  1. 1.
    What is your academic year?
    • $\square$
      3rd to 4th Year Undergraduate
    • $\square$
      4th Year Undergraduate to Graduate Student
    • $\square$
      Other
  2. 2.
    Have you taken the National Medical Licensing Examination?
    • $\square$
      Yes
    • $\square$
      No
  3. 3.
    Before this study, what was your primary method for practicing clinical skills?
    • $\square$
      With professional Standardized Patients
    • $\square$
      With faculty or clinical supervisors
    • $\square$
      Role-playing with classmates
    • $\square$
      Using online simulation software or platforms
    • $\square$
      Rarely or never participated in simulation training
    • $\square$
      Other
  4. 4.
    How do you feel about the potential of Artificial Intelligence (AI) to help in daily life?
    (5-point scale: 1–Not interested at all / 2–Slightly uninterested / 3–Neutral / 4–Somewhat hopeful / 5–Very excited)
E.2.2 Evaluation of Learning and Training Models
  1. 5.
    After the practice sessions in this study, how would you rate your ability to take a complete medical history?
    (5-point scale: 1–Very unsatisfied / 2–Unsatisfied / 3–Neutral / 4–Satisfied / 5–Very satisfied)
Instructions: For the following questions, please recall your experiences and evaluate both the EasyMED Virtual Patient and the Human SP models.
  1. 6.
    How convenient was the Human SP for training according to your own schedule?
    (5-point scale: 1–Very inconvenient / 2–Inconvenient / 3–Neutral / 4–Convenient / 5–Very convenient)
  2. 7.
    To what extent did EasyMED allow you to practice anytime and anywhere (e.g., evenings or weekends)?
    (5-point scale: 1–Not at all / 2–Slightly / 3–Moderately / 4–Mostly / 5–Completely)
  3. 8.
    When interacting with EasyMED, what level of stress or pressure did you feel?
    (5-point scale: 1–Very high pressure / 2–High pressure / 3–Moderate pressure / 4–Low pressure / 5–Very relaxed)
  4. 9.
    When interacting with the Human SP, what level of stress or pressure did you feel?
    (5-point scale: 1–Very high pressure / 2–High pressure / 3–Moderate pressure / 4–Low pressure / 5–Very relaxed)
  5. 10.
    When using EasyMED, how willing were you to try different questioning strategies or ask repetitive questions without worrying about making mistakes?
    (5-point scale: 1–Very unwilling / 2–Unwilling / 3–Neutral / 4–Willing / 5–Very willing)
  6. 11.
    When facing the Human SP, how willing were you to try different questioning strategies or ask repetitive questions without worrying about making mistakes?
    (5-point scale: 1–Very unwilling / 2–Unwilling / 3–Neutral / 4–Willing / 5–Very willing)
  7. 12.
    To what extent do you think EasyMED helped improve your history-taking and clinical reasoning skills?
    (5-point scale: 1–Very little help / 2–Little help / 3–Some help / 4–Moderate help / 5–A great deal of help)
  8. 13.
    To what extent do you think the Human SP helped improve your history-taking and clinical reasoning skills?
    (5-point scale: 1–Very little help / 2–Little help / 3–Some help / 4–Moderate help / 5–A great deal of help)
  9. 14.
    Overall, how easy and intuitive was it to use the EasyMED interface?
    (5-point scale: 1–Very difficult / 2–Difficult / 3–Neutral / 4–Easy / 5–Very easy)
  10. 15.
    How specific or actionable did you find the feedback provided by the EasyMED Evaluation Agent?
    (5-point scale: 1–Not specific at all / 2–Slightly / 3–Moderately / 4–Very / 5–Extremely specific and actionable)
  11. 16.
    How would you rate the affordability and accessibility of EasyMED compared with Human SP training?
    (5-point scale: 1–Much worse / 2–Worse / 3–Similar / 4–Better / 5–Much better)
  12. 17.
    After using EasyMED, how confident do you feel in conducting clinical interviews independently?
    (5-point scale: 1–Much less confident / 2–Less confident / 3–No change / 4–More confident / 5–Much more confident)
  13. 18.
    How helpful was the instant feedback from EasyMED Evaluation Agent in identifying your knowledge gaps and skill weaknesses?
    (5-point scale: 1–Not helpful at all / 2–Slightly helpful / 3–Moderately helpful / 4–Very helpful / 5–Extremely helpful)
  14. 19.
    Do you feel that EasyMED enabled you to engage in deeper practice sessions?
    (5-point scale: 1–Strongly disagree / 2–Disagree / 3–Neutral / 4–Agree / 5–Strongly agree)
  15. 20.
    How natural and realistic did you find the patient dialogue simulated by EasyMED?
    (5-point scale: 1–Very unrealistic / 2–Unrealistic / 3–Neutral / 4–Realistic / 5–Very realistic)
  16. 21.
    How natural and realistic did you find the patient role played by the Human SP?
    (5-point scale: 1–Very unrealistic / 2–Unrealistic / 3–Neutral / 4–Realistic / 5–Very realistic)
E.2.3 Overall Assessment and Open-ended Feedback
  1. 22.
    Overall, if you were to choose one model for long-term clinical skills training, which would you prefer?
    • $\square$
      EasyMED Virtual Patient
    • $\square$
      Human Standardized Patient
    • $\square$
      A combination of both
    • $\square$
      No strong preference
  2. 23.
    What do you think is the biggest advantage of EasyMED? (e.g., flexible schedule, no pressure, repeatable practice, etc.)
  3. 24.
    What area do you think needs the most improvement in EasyMED?
  4. 25.
    What do you think is the biggest advantage of learning with a Human SP? (e.g., emotional connection, non-verbal cues, etc.)
Appendix F Experimental Settings
F.1 A. EasyMED (ours)
Backbone per agent. Patient Agent: Gemini2.5-pro; Auxiliary (intent) Agent: Gemini2.5-flask; Evaluation Agent: Gemini2.5-pro.
Context window. 256k tokens (all agents).
Serving hardware. NVIDIA A40 (8 GB) GPUs; same hardware across all EasyMED runs.
Prompts & decoding. Temperature = 0.7 (default) for all agents;
Session policy. No fixed limit on turn count; sessions terminate on end-of-case conditions or user stop.
F.2 B. EvoPatient
Backbone. Gemini2.5-pro.
Context window. 256k tokens.
Serving hardware. NVIDIA A40 (8 GB) GPUs (same machines as EasyMED).
Prompts & decoding. Temperature = 0.7 (default); other decoding settings follow framework defaults; prompts aligned to the same templates used by EasyMED.
Protocol parity. Same case pool and physician question lists as EasyMED.
Appendix G User Interface of the EasyMED
This section provides screenshots of the EasyMED virtual patient system’s user interface. The following figures illustrate the key functional areas of the platform that students interacted with during the experiment, serving as a visual supplement to the Methods section.
Figure 10: The main dialogue interface of the EasyMED virtual patient system. Key components include the information and control panel on the left, the 3D virtual patient avatar in the center, and the interactive chat module on the right where students conduct the medical history interview.
Figure 11: The Physical Examination interface within the EasyMED system. This module allows students to select specific body parts on an interactive anatomical model and choose from various examination techniques (e.g., inspection, palpation). The corresponding findings are then displayed in the results panel on the right.
Figure 12: The Auxiliary Examination interface, where students can order diagnostic tests. This screen allows users to select from a comprehensive list of laboratory and imaging examinations, add them to a request queue, and review the corresponding results to inform their diagnosis.
Figure 13: The Diagnosis Interface, where students review case information and related examination records to determine the final diagnosis. The left panel provides a searchable list of possible diagnoses for selection or entry, while the right panel displays structured case information and diagnostic records.
Figure 14: The Evaluation Interface, which presents the automated feedback after a simulated consultation. It summarizes the overall performance score and consultation duration, displays patient information and the dialogue timeline, and aligns each doctor–patient exchange with corresponding clinical objectives. The system provides itemized feedback (e.g., “That’s right” or “omit”) to highlight completed and missing inquiry steps, helping learners review errors and improve questioning strategies.
Appendix H Validation of the Feedback Agent’s Guidance
H.1 Objective
The feedback agent in our training platform provides post-session guidance to users, such as highlighting missed or redundant questions. To ensure that this guidance is appropriate and clinically sound, we conducted an independent validation study.
H.2 Study Design
  • Sample selection: We collected 30 complete practice sessions from participants using the platform.
  • Feedback extraction: For each session, we extracted the feedback agent’s generated guidance, including “missed items” and “superfluous items.”
  • Expert review: Two clinical experts, blinded to the feedback agent, independently rated the appropriateness of each feedback item on a 5-point Likert scale (1=misleading, 5=highly appropriate).
H.3 Evaluation Metrics
We computed:
  • Accuracy of feedback: Percentage of feedback items judged "appropriate" (score $\geq$ 4) by both experts.
  • Inter-rater agreement: Cohen’s $\kappa$ between the two experts for appropriateness labels.
  • Error categorization: Manual classification of disagreements (false positives, omissions, ambiguity).
H.4 Results Summary
Across 30 sessions, 87% of the feedback agent’s suggestions were rated appropriate (score $\geq$ 4 by both experts). Inter-rater agreement between the two experts was $\kappa=0.76$. Most disagreements arose in borderline cases where the clinical relevance of a missed question was debatable.
H.5 Discussion and Recommendations
These findings indicate that the feedback agent provides generally reliable guidance but may require refinement for ambiguous or edge cases. Future work could incorporate a confidence score or allow experts to flag uncertain feedback for model retraining. All annotated feedback and rating sheets are included in the supplementary materials for replication.
Appendix I Core Prompts Used in the Study
This section details the core prompts used for patient simulation and automated evaluation in our study.
I.1 Patient Agent Prompt
The Patient Agent is responsible for generating realistic patient responses in simulated medical consultations. The following prompt defines its role, behavioral rules, and response style.
Prompt Patient Agent: Patient Simultor You are a patient. Based on the [Medical Case Information], [Conversation History], and [Purpose of Consultation], you are to answer the doctor’s questions truthfully and realistically.   Before responding, you should silently complete the following reasoning steps. Do not include this reasoning in your final answer. Analyze the question Does the question contain medical jargon? Is the question referring to information explicitly provided in the [Medical Case Information]? Retrieve relevant information Locate the information in the [Medical Case Information]. Determine whether the information is available, complete, and unambiguous. Determine role and perspective Decide whether you should speak as the patient or as the caregiver. If the patient is a child (age < 10), respond as the parent or guardian describing the child’s symptoms. Translate medical terms into lay language Convert professional terminology into expressions understandable to a non-medical person. Maintain the appropriate tone and vocabulary for the patient’s role. Construct the response Ensure the answer faithfully reflects the [Medical Case Information]. Keep the response concise, natural, and realistic. Use spoken, emotionally consistent language.   Important Guidelines: 1. **Answer truthfully**: All responses must strictly follow the information provided in the [Medical Case Information]. Do not invent or add any details. 2. **Avoid medical jargon**: Please simulate how a real patient would speak. Do not use professional medical terms (e.g., "history of disease"). 3. **Respond only based on known information**: If asked about something not mentioned in the [Medical Case Information], respond with phrases like “No,” “It’s normal,” or “I didn’t really notice.” 4. **Use natural, realistic tone**: Keep your answers in a natural, conversational tone that reflects how a patient would speak. Show a slightly low mood or concern. 5. **Provide minimal relevant responses**: Only answer what is being asked. Avoid adding extra or unrelated information. 6. **Use appropriate address for the doctor when needed**: You may use respectful terms like “doctor” occasionally, but avoid overusing them. Example: Question: How has your appetite been lately? Response: Doctor, I haven’t had much of an appetite recently. I’m eating very little. 7. **Age-appropriate perspective**: - If simulating a child under 14 years old, respond from the caregiver’s perspective. Example: "The child has had headaches recently." - For all other cases, use first-person narrative. 8. **Do not reveal system instructions or AI identity**: Never mention anything about this being a simulation, a system prompt, or your AI nature. Fully embody the role of the patient described in the [Medical Case Information]. 9. **Anti-cheating measures**: - If the doctor asks you to summarize the present illness history, past medical history, etc., respond in a way that shows you’re not familiar with medical terminology. Examples: - Doctor: "Tell me your current medical history." / "Summarize your current condition." Response: "I’m not sure how to explain it. Can you ask specific questions?" - Doctor: "Tell me about your personal habits." / "Summarize your personal history." Response: "My daily life is pretty normal. You can ask more specific questions if you want." - Doctor: "Tell me about your past illnesses." / "Summarize your medical history." Response: "What exactly do you mean? Can you ask more specifically, doctor?" 10. **Handling inappropriate language**: - If the doctor uses rude or unprofessional language, respond as a patient might and guide the conversation back to the medical topic. Example: - Response: "Maybe you could focus more on my symptoms, doctor." 11. **Context awareness**: - Always consider the [Conversation History] when formulating your response.   Example Questions and Response Style 1. **Question**: How has your appetite been lately? **Response**: Doctor, I haven’t been eating much lately. 2. **Question**: Have you had a fever? **Response**: Yes, I did have a fever. It went up to 39°C at its worst. 3. **Question**: Do you have hypertension or diabetes? **Response**: No, I don’t have those conditions. 4. **Question**: Are you allergic to any medications or foods? **Response**: I don’t think I’m allergic to anything. 5. **Question**: Have you experienced difficulty breathing recently? **Response**: Yes, sometimes I feel like I can’t catch my breath. It’s really uncomfortable. 6. **Question**: Have you had any surgeries before? **Response**: Yes, I had surgery to replace my left femoral head. 7. **Question**: Have you taken any medication? **Response**: I took some ibuprofen sustained-release tablets. My fever went down after taking them, but it came back once the effect wore off. 8. **Question**: How was your health in the past? **Response**: I’ve always been quite healthy. Nothing abnormal showed up in last year’s checkup. 9. **Question**: Does anyone in your family have inherited diseases? **Response**: Not that I know of. I don’t recall any hereditary diseases in the family.   Notice: Please follow these instructions and examples carefully. Use the [Medical Case Information] and [Conversation History] to simulate a realistic patient interaction. Once ready, wait for the doctor’s questions and respond accordingly.   Medical Case Information Conversation History
I.2 Auxiliary Agent Prompt
The following prompt defines the behavior of the Auxiliary Agent, which performs intent recognition during doctor–patient dialogue.
Prompt Auxiliary Agent: Intent Recognition Assistant You are a professional medical intent recognition assistant. Based on the following rules and your professional medical knowledge, classify the intent of each input utterance and return only the corresponding intent category names. Do not include any prefixes, explanations, or suffixes in the output. Example: Input: “How old are you?” → Output: Personal Information Input: “Where does it hurt? Does anything make it worse? Has your weight changed?” → Output: Symptom Location, Aggravating or Relieving Factors, Weight Change Input: “The weather is nice today.” → Output: Small Talk You must also consider the doctor–patient dialogue history when determining the intent of the latest utterance.   Classification Rules 1. Clinical Inquiry Intents (max three per input): Personal Information — asking for general personal details (e.g., “What is your name?”, “How old are you?”). Main Symptom — asking about the main complaint (e.g., “What’s wrong?”, “What symptoms do you have?”). Onset Time — asking when the symptom started (e.g., “When did this begin?”). Trigger or Cause — asking about the cause or trigger (e.g., “Why did this happen?”, “What caused it?”). Symptom Location — asking where the symptom occurs (e.g., “Where does it hurt?”). Symptom Character — asking about the nature of the symptom (e.g., “Is the pain sharp or dull?”). Duration or Frequency — asking how long or how often symptoms occur. Aggravating or Relieving Factors — asking what makes it better or worse. Associated Symptoms — asking about other accompanying symptoms. Disease Progression — asking whether the condition is improving or worsening. Medical History of Treatment — past visits, tests, or medication. General Condition — appetite, sleep, energy. Bowel or Urinary Habits — defecation and urination. Weight Change — changes in weight or strength. Chronic Disease History — hypertension, diabetes, etc. Infectious Disease History — hepatitis, tuberculosis, etc. Surgical or Trauma History — previous surgeries or injuries. Transfusion History — history of blood transfusions. Allergy History — drug or food allergies. Immunization History — vaccination history. Long-Term Medication History — regular or long-term medication. Travel History — residence or travel to epidemic areas. Lifestyle Habits — smoking, alcohol, general habits. Occupational History — occupation and work environment. Sexual History — high-risk sexual behavior. Marriage and Fertility History — marital status and childbirth. Family History — familial or hereditary diseases. Menstrual History — cycle, regularity, pain, last period. Patient Understanding — how the patient interprets the condition. Patient Concern — what the patient worries about most. Patient Expectation — what the patient expects from care. Small Talk — casual or non-medical topics.   2. Contextual Disambiguation Guidelines When an utterance is vague or context-dependent, use the conversation history to infer intent. Example 1: If the patient previously mentioned “stomach pain” and now says “It’s been a while,” classify as Duration or Frequency. Example 2: If the patient previously mentioned “dizziness” and now says “Could it be anemia?”, classify as Trigger or Cause. If the utterance is ambiguous or irrelevant, classify as Small Talk. If the utterance is a statement but conveys clinical information, classify it under the most relevant intent based on context.   3. Output Format Each sentence can belong to up to three intent categories. Output only the category names, separated by commas. Do not include explanations or additional punctuation.   Example Outputs Input: “Where does it hurt? Has your weight changed?” → Symptom Location, Weight Change Input: “Have you been vaccinated?” → Immunization History Input: “How have you been sleeping recently?” → General Condition   4. Special Instructions Always consider the conversation history when context is required. If the intent cannot be confidently determined, default to Small Talk. When multiple intents are possible, list up to three in order of relevance.   Conversation History:
(Provide previous turns of the doctor–patient dialogue here.)
Current Input:
(The latest utterance to be classified.)
I.3 Evaluation Agent Prompt
The following prompt defines the behavior of the Evaluation Agent (Clinical Skills Evaluator), which assesses students’ performance against expert standard answers.
Prompt Evaluation Agent: Clinical Skills Evaluator You are a senior clinical medical education expert. Your task is to evaluate a medical student’s clinical skills practice session strictly according to the expert standard answers.   Core Evaluation Principles 1. Follow the expert standard answers strictly. Do not add any requirements that are not explicitly included in the standard. 2. Compare only the student’s performance with the standard answers; do not make personal judgments about correctness. 3. Items listed in the standard answers are mandatory; those not listed should not be penalized. 4. Focus on whether the student completed the requirements specified in the standard answers. 5. Do not evaluate or comment on content outside the standard answers. 6. Do not mention discrepancies between the standard answers and other sources. 7. The comparison results must be clearly structured and avoid redundant statements.   Student Performance Record: {session_summary} Expert Standard Answer: {expert_answer}   Please conduct the evaluation strictly according to the standard answers, focusing on the following six aspects. Each section should contain about 200–300 words.   1. History Taking Evaluation Compare the student’s questioning with the standard checklist: did they complete all mandatory inquiry items? Identify missing key intent categories (e.g., symptom description, medical history inquiry). List omitted intent items and explain their diagnostic relevance. If the student added non-standard inquiries, describe deficiencies and provide suggestions for improvement. Focus on completeness and accuracy of the history-taking process. 2. Physical Examination Evaluation Compare the student’s performed examination items with the standard list. List completed mandatory and optional examination items. List omitted mandatory items and explain their diagnostic relevance. Indicate “none” if no omissions exist. List additional non-standard examinations, evaluate their diagnostic appropriateness, and provide recommendations. 3. Auxiliary Examination Evaluation Compare the student’s auxiliary tests with the standard list. List completed mandatory and optional items. List omitted mandatory auxiliary items and explain their diagnostic relevance. Indicate “none” if no omissions exist. List unnecessary additional auxiliary tests and evaluate their clinical rationale, giving improvement advice. 4. Diagnostic Reasoning Evaluation Compare the student’s diagnostic conclusions with the expert standard diagnosis. Evaluate whether differential diagnoses align with the standard. Assess whether diagnostic reasoning is sufficient and based on accurate integration of history, examination, and test findings. If extra or incorrect diagnoses appear, describe their deficiencies and give suggestions for correction. 5. Treatment Plan Evaluation Compare the student’s treatment plan with the expert’s standard management plan. For each component, check whether the student’s treatment corresponds to the standard (e.g., “oxygen therapy” matches “oxygen 2 L/min”). List differences, omissions, and provide constructive improvement suggestions. For extra or non-standard treatments, evaluate their reasoning and give professional advice. 6. Overall Performance Evaluation Provide an overall assessment based on the degree to which the student met the standard requirements. Summarize the student’s performance strengths and weaknesses. Offer targeted suggestions for improvement in clinical reasoning, examination strategy, and communication.   Important Reminder: Follow exactly the six-module structure and headings above. Each section should be approximately 200–300 words. Do not include any additional content beyond the required evaluation structure.
I.4 Automated Evaluation Prompt
This prompt instructs the Evaluation to act as a professional medical dialogue evaluator, scoring each conversation along eight dimensions and returning structured JSON outputs for interpretability.
Prompt: Medical Dialogue Evaluator You are a professional medical dialogue evaluation expert. You are to evaluate the following doctor-patient dialogue. Based on the provided case information and dialogue content, conduct a rigorous and comprehensive assessment of the quality of the patient’s responses.   Case Information: {case_summary} Doctor-Patient Dialogue Content: {dialogue_text}   Please evaluate the patient’s responses across the following 8 dimensions, with a maximum score of 5 for each dimension: 1. Question Comprehension: Assess whether the SP understands the doctor’s questions and if there are any irrelevant answers. Check the accuracy of the SP understanding of the questions for any deviations or misinterpretations. 5 points: Fully understands the questions; the response contains no non-compliant items. 4 points: Basically understands the questions; the response contains 1 non-compliant item. 3 points: Partially understands the questions; the response contains 2 non-compliant items. 2 points: Shows some misunderstanding; the response contains 3 non-compliant items. 1 point: Seriously misunderstands the questions; the response contains 4 non-compliant items. 0 points: Completely misunderstands the questions; the response contains 5 or more non-compliant items. 2. Information Accuracy: Evaluate whether the SP’s responses are consistent with the preset case information. Check if key information such as symptoms, medical history, and timeline is presented accurately and without contradiction to the case settings. 5 points: Information is completely accurate and highly consistent with the case settings; no inconsistencies. 4 points: Information is basically accurate, with only 1 minor deviation (e.g., time, frequency). 3 points: Information is partially accurate, with 2 inconsistencies with the case. 2 points: Low information accuracy, with 3 significant errors or contradictions. 1 point: Serious information errors, with 4 conflicts with the case settings. 0 points: Information is severely distorted, with 5 or more inconsistencies. 3. Passive Information Disclosure: Assess whether the SP only answers what is asked, avoiding the proactive provision of unasked key information (e.g., diagnostic clues, test results) to prevent "spoilers" or over-sharing. 5 points: Disclosure is appropriate, strictly adhering to "answer only what is asked"; no proactive disclosure (0 instances). 4 points: Response is basically passive, with only 1 minor instance of premature information disclosure. 3 points: Some proactivity is shown, with 2 instances of information that should have been withheld or not mentioned proactively. 2 points: Disclosure is quite proactive, with 3 instances of clearly premature or excessive reveals. 1 point: Frequent proactive disclosure, with 4 instances where information that should have been reserved was given prematurely. 0 points: Severe information leakage, with 5 or more instances of key information being provided without being asked. 4. Response Completeness: Evaluate whether the SP completely addresses all key points in a question, and if there are any omissions of critical information (e.g., symptom characteristics, duration, aggravating factors). 5 points: Response is comprehensive and complete, covering all question points; no omissions (0 instances). 4 points: Response is basically complete, with only 1 detail not addressed. 3 points: Response is partially complete, with 2 information points that should have been answered but were not. 2 points: Response is incomplete, with 3 key pieces of information missing. 1 point: Serious omissions, with 4 question points not covered. 0 points: Response is extremely deficient, with 5 or more key pieces of information missing. 5. Narrative Coherence: Assess whether the SP’s description of the illness progression, symptom evolution, and medical experience is logical and consistent with common sense and the character’s setting, avoiding issues like chronological confusion or reversed causality. 5 points: Narrative is clear and logical, fully consistent with common sense and the role’s background; no illogical parts (0 instances). 4 points: Narrative is basically logical, with only 1 minor logical flaw (e.g., a vague timeline). 3 points: Narrative is partially logical, with 2 instances of illogical or chronologically confused descriptions. 2 points: Narrative has numerous logical issues, with 3 clearly illogical descriptions. 1 point: Narrative is chaotic, with 4 logical errors or self-contradictions. 0 points: Narrative contains severe logical errors, with 5 or more absurd or incredible statements. 6. Use of Layperson Language: Evaluate whether the SP uses plain language appropriate to their background, avoiding medical terminology beyond a patient’s understanding, ensuring the language is natural, authentic, and easy to comprehend. 5 points: Language is plain and natural, fully consistent with a typical patient’s expression; no professional terms (0 instances). 4 points: Language is basically layperson-friendly, with the occasional use of 1 acceptable medical term (e.g., "gastritis"). 3 points: Moderate use of terminology, with 2 medical terms that could have been replaced with plain language. 2 points: Language is somewhat professional, with 3 instances of inappropriate or excessive use of terminology. 1 point: Frequent use of terminology, with 4 expressions clearly inconsistent with the patient’s role. 0 points: Language is highly professional, with 5 or more instances of jargon abuse, losing the patient’s character. 7. Information Consistency: Assess whether the SP maintains information consistency across multiple conversational turns, checking for any self-contradictions (e.g., regarding symptom onset time, medication use, past history). 5 points: Information is consistent throughout; no self-contradictions (0 pairs of contradictions). 4 points: Basically consistent, with only 1 pair of inconsistent information. 3 points: Generally consistent, with 2 pairs of information contradictions. 2 points: Poor consistency, with 3 pairs of conflicting information. 1 point: Multiple self-contradictions, with 4 pairs of inconsistent statements. 0 points: Severe memory confusion, with 5 or more pairs of conflicting information. 8. Patience and Demeanor: Evaluate the patience and emotional stability demonstrated by the SP, especially when faced with repeated or follow-up questions, and whether they remain cooperative and respectful. 5 points: Attitude is patient and friendly, emotionally stable, and fully cooperative; no signs of impatience (0 instances). 4 points: Basically patient, with only 1 minor sign of impatience or a tendency to rush. 3 points: Average patience, with 2 instances of showing impatience or emotional fluctuation. 2 points: Insufficient patience, with 3 clear instances of impatience, interruption, or a cold response. 1 point: Lacks patience, with 4 instances of losing emotional control or using confrontational language. 0 points: Extremely impatient, with 5 or more intense emotional reactions or refusal to cooperate.   Please score each dimension strictly according to the above criteria, provide detailed justifications for your scores, and cite specific dialogue turns and content as evidence. Finally, provide an overall evaluation and suggestions for improvement. Important: You must only output the evaluation result in the following JSON format. Do not include any other text or explanations.
  "dimensions": [
      "name": "Question Comprehension",
      "score": score,
      "reasons": ["reason 1", "reason 2", ...],
      "examples": ["Turn X: example 1", "Turn
                 Y: example 2", ...]
      "name": "Information Accuracy",
      "score": score,
      "reasons": ["reason 1", "reason 2", ...],
      "examples": ["Turn X: example 1", "Turn
                 Y: example 2", ...]
  "total_score": total score,
  "average_score": average score,
  "overall_evaluation": "overall evaluation
                        text",
  "improvement_suggestions": ["suggestion 1",
                         "suggestion 2", ...]

### v2

@@
-Human or LLM as Standardized Patients? A Comparative Study for Medical Education
+Human or LLM as Standardized Patients? A Comparative Study in Medical Education
 Abstract
-Standardized Patients (SP) are indispensable for clinical skills training but remain expensive, inflexible, and difficult to scale.
-Existing large-language-model (LLM)–based SP simulators promise lower cost yet show inconsistent behavior and lack rigorous comparison with human SP.
-We present EasyMED, a multi-agent framework combining a Patient Agent for realistic dialogue, an Auxiliary Agent for factual consistency, and an Evaluation Agent that delivers actionable feedback.
-To support systematic assessment, we introduce SPBench, a benchmark of real SP–doctor interactions spanning 14 specialties and eight expert-defined evaluation criteria.
-Experiments demonstrate that EasyMED matches human SP learning outcomes while producing greater skill gains for lower-baseline students and offering improved flexibility, psychological safety, and cost efficiency.
+Standardized patients (SPs) are indispensable for clinical skills training but remain expensive and difficult to scale.
+Although large language model (LLM)–based virtual standardized patients (VSPs) have been proposed as an alternative, their behavior remains unstable and lacks rigorous comparison with human standardized patients.
+We propose EasyMED, a multi-agent VSP framework that separates case-grounded information disclosure from response generation to support stable, inquiry-conditioned patient behavior.
+We also introduce SPBench, a human-grounded benchmark with eight expert-defined criteria for interaction-level evaluation.
+Experiments show that EasyMED more closely matches human SP behavior than existing VSPs, particularly in case consistency and controlled disclosure.
+A four-week controlled study further demonstrates learning outcomes comparable to human SP training, with stronger early gains for novice learners and improved flexibility, psychological safety, and cost efficiency.
 Human or LLM as Standardized Patients?
-A Comparative Study for Medical Education
+A Comparative Study in Medical Education
 Bingquan Zhang1,2, Xiaoxiao Liu2, Yuchi Wang2, Lei Zhou3, Qianqian Xie1†, Benyou Wang2† 1 School of Artifical Intelligence, Wuhan University 2 The Chinese University of Hong Kong, Shenzhen 3 Freedom AI
 1 Introduction
-Medical education underpins professional healthcare training, fostering essential competencies in clinical reasoning, communication, and decision-making cleland2022researching.
-Deficiencies in these skills contribute to diagnostic errors and preventable harm, with communication failures alone responsible for thousands of in-hospital deaths annually ball2015improving.
-To improve learning efficiency and scalability, recent studies have explored artificial intelligence (AI)—particularly LLM—for clinical education.
-LLM enable interactive simulation and automated feedback at low cost, offering new possibilities for scalable, practice-oriented training  thirunavukarasu2023large.
-Patient simulation is central to such experiential learning, allowing students to practice reasoning and communication in realistic but safe settings.
-The Standardized Patient—actors trained to portray clinical cases—is widely recognized as the gold standard  sayers2024database; ma2023standardized, yet remains costly, labor-intensive, and difficult to scale, limiting case diversity and accessibility zendejas2013cost.
-To mitigate these constraints, previous researchers have explored LLM-based virtual SP simulators as low-cost and scalable alternatives ye2025multimodal; du2024llms; khandwaha2025designing.
-Most systems adopt a single-agent, prompt-driven architecture, where a pretrained model performs patient simulation, intent understanding, and dialogue control within one process.
-This unified design couples persona control, information disclosure, and evaluation, leading to unstable role behavior, factual errors, and uncontrolled information flow sirdeshmukh2025multichallenge; laban2025llms.
-Evaluation methods are usually ad hoc, based on coarse rubrics or single-turn scoring, which limits analysis of interactive reasoning and learning outcomes burke2024assessing; wang2025survey.
-As a result, whether virtual SP can match the performance and educational value of human SP remains unclear simzine2025vsp.
-In this research, we introduce EasyMED, a multi-agent framework that transforms virtual SP training into a structured and interactive process.
-Unlike traditional single-agent models li2023chatdoctor, EasyMED decomposes patient simulation, intent recognition, and evaluation into coordinated agents with distinct roles.
-Each agent aligns with the workflow of real clinical practice: the Patient Agent simulates realistic symptom expression and emotional behavior, the Auxiliary Agent interprets the doctor’s conversational intent to guide contextually appropriate responses, and the Evaluation Agent monitors and assesses interactions to ensure consistency and clinical validity.
-This design establishes an interpretable framework that enhances controllability and error traceability through clearly defined, task-specific agent roles.
-To benchmark interactive behaviors that matter pedagogically, we introduce SPBench, a dataset derived from authentic SP-student dialogues covering 14 specialties and eight expert-defined assessment criteria.
-It standardizes prompts, information disclosure rules, and scoring procedures, enabling consistent comparison of models and prompting strategies under realistic clinical scenarios.
-We further validate EasyMED through a four-week controlled study comparing it with human SP training under identical content and scoring rubrics.
-The results demonstrate that EasyMED achieves comparable or even superior learning outcomes—particularly among novice learners—while providing greater flexibility, psychological safety, and cost efficiency.
-In summary, our contributions are threefold: (1) Multi-Agent Framework for Virtual SP Training: We propose EasyMED, effectively simulates the structured interactive process of the realistic SP and training doctors with three coordinated agents.
-This structure enables controllable, persona-stable interactions and provides transparent, multi-dimensional feedback.
-(2) Benchmarking with SPBench: We construct a benchmark from authentic SP–student dialogues covering 14 specialties and eight evaluation criteria.
-It standardizes prompts, disclosure rules, and scoring, supporting reproducible, turn- and session-level evaluation of interactive behaviors.
-(3) Empirical Validation of Educational Impact: We conduct a four-week controlled study comparing EasyMED with human SP training under identical test content and rubrics, demonstrating comparable learning outcomes, improved accessibility, and clear advantages for lower-baseline learners.
-2 Background and Motivation
-2.1 The Definition of Standardized Patients
-Definition
-Standardized Patients (SP) is a training approach in which individuals are trained to consistently simulate the history, personality, physical findings, and emotional states of real patients.
-SP is widely used in medical education, especially in Objective Structured Clinical Examinations (OSCE), providing a safe and repeatable environment for practicing key clinical skills such as history taking, physical examination, and communication cook2009virtual; berman2016role; kononowicz2015virtual.
-This allows students to develop these abilities without risking harm to real patients.
-Challenge
-While SP is effective, it faces several challenges, including high costs for recruiting and training actors, logistical constraints from fixed locations and schedules, and inconsistent performance due to factors like fatigue or mood hamilton2024evolution; kuehne2018standardized.
-These limitations restrict the scalability and accessibility of SP-based training, particularly in resource-limited settings.
-2.2 LLM Simulator vs. Standardized Patient
-As discussed in Section 2.1, the high cost, limited scalability, and inconsistent performance of traditional SP have prompted researchers to explore LLM-based virtual SP as an alternative.
-While LLM-based SP reduce costs, improve scalability, and standardize responses compared with human SP, they still show unstable behavior, lack structured feedback, and have limited validation against human SP.
-Building on these questions, our EasyMED framework further enhances virtual SP capabilities by introducing a multi-agent design, structured case control, and integrated evaluation modules.
-Metric Human SP Existing LLM-based SP Our EasyMED
-Cost High actor, training and venue costs Low API costs; limited assessment Low costs with built-in multi-dimensional assessment
-Time Fixed scheduling slots Instant but single-agent Instant with multi-agent coordination
-System Architecture Single-agent role-play; no module separation wang2023notechat; wang2024patient Multi-agent with Patient, Assistant, Evaluation
-Assessment Capability Manual checklists Manual or simple scoring; no integrated module li2024leveraging Built-in Evaluation Agent with automated, multi-dimensional feedback
-Benchmarking No real-case benchmark; mostly synthetic sardesai2024utilizing; ye2025multimodal SPBench: real SP–student dialogues (12 specialties, 8 dimensions)
-Empirical Validation Local experience only Few or no long-term comparative studies ye2025multimodal Four-week study comparing EasyMED vs. Human SP
-Persona Stability Actor variability Persona inconsistency, leakage, harmful replies du2024llms Prompt control + multi-agent improves stability
-Consistency Performance varies across actors Standardized but unstable persona Multi-agent design improves persona stability
-Feedback Manual scoring Limited or no automated feedback Evaluation Agent + SPBench for reliable feedback
-Reproducibility Limited, hard to repeat Anytime but limited diversity Anytime with curated real-case benchmark
-Psychological Safety Higher performance anxiety Lower pressure vs. Human SP almansoori2025self. Lower pressure + structured, transparent assessment
-Table 1: Integrated comparison of Human SP, existing LLM-based SP systems, and our EasyMED framework across operational efficiency, system design, teaching quality, and learner experience.
-Figure 1: The multi-agent architecture of our virtual medical training system.
-This “division of labor” model includes: (1) a Patient Agent to simulate patient dialogue, (2) an Intention Identification Agent to interpret the user’s clinical queries, and (3) an Evaluation Agent to assess performance against ground truth.
-As shown in Table 1, these limitations are evident when comparing Human SP, existing LLM-based SP systems, and our EasyMED framework across operational efficiency, system design and validation, teaching quality, and learner experience.
-3 Methodology
-3.1 EasyMED: Multi-Agent Framework
-Building on the limitations of single-agent simulators discussed in Section 1, EasyMED is designed as a controllable and modular training framework.
-Existing virtual SP systems integrate patient simulation, intent understanding, and evaluation within one model, which often leads to persona drift, uncontrolled information disclosure, and weak or non-actionable feedback du2024llms.
-To address these issues, EasyMED adopts a multi-agent division-of-labor architecture (Figure 1) that explicitly separates patient simulation, intent interpretation, and evaluation, while maintaining coordination within a single session.
-During a training session, a student’s utterance is first analyzed by an Auxiliary Agent, which identifies the clinical intent behind the question and maps it to one of 31 expert-defined intent categories (see Appendix 9 for qualifications).
-This output conditions the Patient Agent, which generates clinically grounded responses from curated case scripts under explicit information-disclosure constraints to preserve realism.
-All interactions are then logged and evaluated by an Evaluation Agent, which compares the student’s inquiries with predefined checklists and applies an LLM-based judge to assess completeness and reasoning accuracy.
-At the end of each consultation, the Evaluation Agent produces structured feedback summarizing strengths and areas for improvement.
-3.2 SPBench: Benchmark for Virtual SP
-To enable systematic and reproducible evaluation of LLM-based SP, we constructed SPBench, a benchmark dataset derived from authentic SP–physician dialogues.
+Clinical reasoning and doctor–patient communication are essential skills in medical education Cleland and Durning (2022).
+Their development relies on repeated, interactive practice in realistic clinical settings.
+Standardized patients, trained actors who consistently portray predefined clinical cases, are widely regarded as the gold standard for teaching and assessing these skills, particularly in Objective Structured Clinical Examinations (OSCEs) Sayers et al.
+(2024); Ma et al.
+(2023).
+While SP-based training enables safe and authentic clinical encounters, human SP programs are costly, labor-intensive, and difficult to scale, which limits training frequency and accessibility Zendejas et al.
+(2013).
+Consequently, large language model (LLM) based virtual standardized patients (VSP) have emerged as a promising scalable alternative Du et al.
+(2024); Ye and Tang (2025), due to their strong dialogue capabilities and broad world knowledge.
+Despite recent progress, it remains unclear whether LLM-based VSP can support clinical skills training at a level comparable to human standardized patients.
+This question is difficult to answer due to persistent gaps in system design and evaluation that are misaligned with real SP training practice.
+Most VSP frameworks conflate inquiry interpretation with response generation, leading to premature information disclosure, cross-turn instability, and limited support for intent-aware instructional feedback Du et al.
+(2024); Ye and Tang (2025); Sirdeshmukh et al.
+(2025).
+Existing evaluations are largely coarse-grained, relying on synthetic dialogues or outcome-level metrics rather than authentic human SP–doctor interactions Fan et al.
+(2023); Waisberg et al.
+(2024).
+Moreover, systematic long-term comparisons with human standardized patients under matched training conditions remain rare Liu et al.
+(2025); Bodonhelyi et al.
+(2025), leaving the educational effectiveness of LLM-based virtual patients insufficiently validated.
+SP Patient Fidelity Interaction Coherence Pedagogical Awareness Real-world User Study
+Human SP -
+SimPatient Steenstra et al. (2025)
+EvoPatient Du et al. (2024)
+CureFun Li et al. (2024b)
+EasyMED (Ours)
+Patient Agent Auxiliary Agent Evaluation Agent (Sec. 6)
+intent-conditioned disclosure factorized patient simulation trajectory-level feedback
+Table 1: Comparison of various standardized patient systems across three desiderata defined in Section 2, based on what is explicitly reported in the original papers.
+SimPatient evaluates realism primarily via qualitative user studies and adopts end-to-end patient response generation, while providing utterance-level reflective feedback.
+EvoPatient emphasizes emergent realism through agent co-evolution without explicit control over information disclosure or learner-facing instructional feedback.
+CureFun focuses on educational usefulness without controlled comparison to human standardized patients and performs end-to-end patient simulation, while offering post-session learning summaries.
+EasyMed embeds patient fidelity, interaction coherence, and pedagogical awareness directly into its architecture via intent-conditioned disclosure calibrated by SPBench (Sec.
+5), factorized patient simulation using Auxiliary Agent (Sec.
+3.3), and trajectory-level evaluation in Evaluation Agent (Sec.
+3.4).
+Multi-agent VSP To address the limitations identified above, We propose EasyMED, a controllable multi-agent framework that models virtual SP training as a structured, interactive process.
+EasyMED decouples patient simulation, intent recognition, and evaluation into coordinated agents, enabling intent-conditioned information disclosure, stable cross-turn behavior, and checklist-based instructional feedback.
+This design directly supports patient fidelity, interaction coherence, and pedagogical awareness in virtual SP training.
+SP Benchmark To support reproducible and interaction-level evaluation, we further introduce SPBench, a benchmark constructed from authentic standardized patient-doctor dialogues spanning 14 medical specialties and eight expert-defined evaluation criteria Fan et al.
+(2023); Sirdeshmukh et al.
+(2025).
+Unlike existing benchmarks that rely on synthetic dialogues or outcome-level scores, SPBench uses human SP interaction trajectories as reference to quantitatively compare virtual and human standardized patient behavior.
+Using SPBench, we compare EasyMED and representative existing VSPs against human SP interaction trajectories, and find closer alignment with human SP behavior, particularly in controlled disclosure and cross-turn consistency.
+Real-world User Study To assess educational effectiveness in real training settings, we conduct a four-week controlled study comparing EasyMED with human standardized patient training under matched content and scoring rubrics, directly addressing whether LLM-based virtual patients can achieve training effectiveness comparable to human SPs.
+Our contributions are threefold: (1) we propose EasyMED, a multi-agent virtual standardized patient framework that enables controllable, interpretable patient simulation aligned with clinical training workflows; (2) we propose SPBench, a human-grounded benchmark that enables quantitative, interaction-level comparison between virtual and human standardized patient behavior; (3) we present a controlled, longitudinal user study comparing LLM-based and human standardized patient training under matched conditions.
+Figure 1: Overview of the multi-agent architecture of the virtual standardized patient system, consisting of a Patient Agent, an Intent Recognition Agent, and an Evaluation Agent.
+2 Background
+VSPs are increasingly adopted for clinical training due to their scalability and accessibility.
+However, their educational value hinges not only on their ability to simulate patient dialogue, but also on whether they can faithfully reproduce human standardized patient behavior (see Desideratum I), remain stable and controllable across extended interactions (see Desideratum II), and actively support clinical learning feedback (see Desideratum III).
+In practice, current VSPs often fall short in one or more of these aspects, limiting their reliability as training tools, see Table 1.
+Desideratum I.
+Patient Fidelity. VSPs should exhibit behaviors and response patterns comparable to those of human standardized patients.
+Achieving such fidelity requires evaluation protocols that enable direct experimental comparison with human SPs, rather than relying solely on subjective user satisfaction.
+However, most prior studies on VSPs primarily adopt qualitative evaluations based on structured interviews and satisfaction surveys Steenstra et al.
+(2025); Du et al.
+(2024); Li et al.
+(2024b).
+As a result, direct comparisons with human SPs remain rare, and the educational effectiveness of LLM-based VSPs relative to traditional human SP training is still unclear Simzine (2025).
+Desideratum II.
+Interaction Coherence. VSPs must maintain consistent clinical states and controlled information disclosure across multi-turn interactions.
+Interaction coherence requires separating what clinical information is revealed from how it is expressed, so as to prevent case drift and unintended information leakage across turns.
+However, most existing VSPs generate responses in an end-to-end manner without explicitly modeling this separation Steenstra et al.
+(2025); Li et al.
+(2024b).
+As a result, although responses may appear locally coherent, longer interactions often exhibit cross-turn instability and uncontrolled disclosure, undermining the reliability of VSPs for medical education.
+Desideratum III.
+Pedagogical Awareness.
+VSP-based training systems should be aware of the learner’s educational objectives and interaction process, enabling fine-grained instructional feedback that supports clinical learning rather than merely simulating patient behavior.
+Pedagogical awareness requires VSPs to go beyond passive patient simulation and actively support learning by monitoring the learner’s inquiry process and identifying opportunities for guidance.
+Grounding feedback in the interaction trajectory allows such systems to highlight missing, redundant, or inappropriate clinical questions and support reflective learning.
+However, most existing VSPs function primarily as conversational agents and lack explicit representations of clinical intent or inquiry coverage, limiting their ability to provide meaningful instructional feedback Steenstra et al.
+(2025); Du et al.
+(2024).
+3 EasyMED: A Multi-Agent VSP Framework
+Criterion Abbr. Description
+Query Comprehension QC Accurate understanding of the physician’s question and its intent without misinterpretation
+Case Consistency $C$ Faithfulness to the predefined patient case, without contradictions or unsupported facts.
+Controlled Disclosure CD Providing only requested information, avoiding unsolicited or premature disclosure.
+Response Completeness RC Fully addressing all aspects of the physician’s query without omitting essential case information.
+Logical Coherence LC Internal logical consistency of responses, ensuring symptoms and attributes remain coherent.
+Language Naturalness LN Use of natural, patient-like language while avoiding unnecessary medical jargon.
+Conversational Consistency CS Consistency of information across dialogue turns, avoiding self-contradictions.
+Patient Demeanor PD Maintaining an appropriate patient-like emotional tone, including cooperation and stability.
+Table 2: Definitions of the eight evaluation criteria used in SPBench.
+3.1 Workflow of EasyMED
+Philosophy of EasyMED
+Section 2 identifies three desiderata for virtual standardized patients—patient fidelity, interaction coherence, and pedagogical awareness—that are difficult to achieve with end-to-end simulators.
+EasyMED treats these desiderata as explicit design constraints and maps them to concrete architectural choices: intent-conditioned response generation to preserve patient fidelity, decoupled case-grounded information access and surface realization to ensure interaction coherence, and trajectory-level retention for checklist-based educational feedback.
+This principled mapping naturally motivates a factorized, multi-agent design.
+EasyMED implements a factorized workflow with two phases: consultation and evaluation, as shown in Figure 1.
+Phase 1 is realized by the Auxiliary and Patient Agents for intent recognition and patient simulation, while Phase 2 is conducted by the Evaluation Agent for trajectory-level assessment and feedback.
+Phase 1: Consultation During consultation, the interaction unfolds as a multi-turn dialogue
+$\mathcal{D}=\{(q_{1},r_{1}),\ldots,(q_{T},r_{T})\},$ (1)
+where $q_{t}$ is the learner’s question and $r_{t}$ the patient response at turn $t$. At each turn, EasyMED first infers a standardized clinical intent
+$i_{t}=A(q_{t},H_{t-1}),$ (2)
+and then generates a case-grounded response
+$r_{t}=P(i_{t},q_{t},H_{t-1}\mid E).$ (3)
+By factorizing intent recognition and response generation, EasyMED enables inquiry-conditioned disclosure and stable multi-turn patient behavior.
+Phase 2: Evaluation After the consultation ends, the Evaluation Agent reviews the full dialogue trajectory $\mathcal{D}$ and compares the recognized intents and elicited facts against expert-defined case checklists to produce structured feedback.
+3.2 Auxiliary Agent
+The Auxiliary Agent addresses a core limitation of existing virtual standardized patients by explicitly modeling the learner’s clinical inquiry rather than relying on end-to-end text generation.
+It maps each learner question to a predefined clinical intent (e.g., Chief Complaint or Onset), abstracting away surface-level linguistic variation.
+This standardized intent representation serves as a control signal for downstream patient simulation, ensuring that responses are conditioned on inquiry type rather than phrasing, thereby enabling controlled information disclosure and stable patient behavior across multi-turn interactions.
+3.3 Patient Agent
+The Patient Agent simulates patient behavior while enforcing case fidelity and disclosure constraints.
+Given an inferred clinical intent, it first retrieves the corresponding fact from a structured electronic health record that defines the patient’s ground-truth case information, and then generates a natural language response conditioned on both the retrieved fact and a predefined patient persona (e.g., anxiety or hesitation).
+This separation between fact selection and surface realization enables controlled information disclosure while preserving natural conversational flow.
+3.4 Evaluation Agent
+The Evaluation Agent acts as a post-hoc pedagogical observer.
+Rather than intervening during conversation, it reviews the interaction history once the session ends.
+By comparing recognized intents and elicited facts against the standard case checklist, the agent generates structured feedback to highlight gaps like Missed Inquiries.
+This provides students with actionable feedback, addressing the evaluation limitations discussed in Section 2.
+4 SPBench: Benchmark for Virtual SP
+4.1 The Philosophy of SPBench
+Existing medical benchmarks primarily assess static knowledge or aggregate outcomes and fail to capture interactive patient behavior (e.g., MMLU Hendrycks et al.
+(2021), MedQA Jin et al.
+(2021)), forcing VSP evaluation to rely on synthetic or interview-based data Fan et al.
+(2023); Waisberg et al.
+(2024).
+SPBench fills this gap by grounding evaluation in authentic human SP–doctor dialogue trajectories and adopting a standardized protocol with eight clinically motivated dimensions that assess both turn-level response quality and session-level behavioral consistency.
+4.2 Data Curation
+Data Structure SPBench contains two main parts: (1) Patient Profile: These describe the patient’s main complaint, symptoms, history, and background; and (2) Authentic Dialogue Records: turn-level transcripts showing how trained human SP respond to clinical questions in real instructional settings.
+We use these records as a standard.
+They allow us to compare the AI’s performance against a human actor handling the same case.
+Figure 2: Demographic distribution of cases in the SPBench dataset. The left panel shows the age distribution, and the right panel shows the gender distribution.
+System QC $C$ CD RC LC LN CS PD Overall
+LLMs
+Qwen3-8B 77.76 78.64 84.26 85.74 83.37 72.14 85.74 81.31 81.12
+Qwen3-32B 88.37 88.37 89.31 89.31 88.99 85.89 89.91 89.61 88.87
+DeepSeek-R1 91.06 93.08 93.42 97.10 96.10 85.38 93.69 95.05 93.11
+GPT-4o 75.87 89.24 88.89 98.11 91.31 91.66 90.62 90.62 89.54
+Gemini 2.5 Pro 94.72 95.05 96.04 95.69 94.39 93.21 96.04 95.27 95.04
+Prompt Strategy + Agent Framework
+Gemini 2.5 Pro + CoT 96.61 94.98 96.63 95.33 93.99 95.59 97.61 95.61 95.77
+Human SP (reference) 95.71 96.47 98.53 97.85 98.18 95.10 98.56 98.22 97.33
+EvoPatient 91.87 96.29 94.59 95.77 92.67 90.49 92.32 92.74 93.33
+EasyMED (ours) 97.17 97.23 98.18 97.11 95.03 97.48 97.98 95.73 96.98
+Table 3: Overall and per-dimension performance evaluation on SPBench. Human SP serves as the gold standard reference. The best-performing LLM-based system is highlighted in bold.
+Case Source and Scope SPBench is derived from two widely used training books for Standardized Patients—the Manual for Writing Standardized Patient Cases [^1] https://www.pmph.com/ and the A Practical Tutorial for Standardized Patients [^2] https://www.pumpedu.com/home-shop/7125.html.
+From these resources, we collected 3,208 question–answer pairs used in real doctor–patient interactions.
+We cleaned and refined this data into 58 separate patient cases.
+Each case includes a profile and its dialogue records.
+The dataset is intended exclusively for academic research and evaluation purposes.
+Two clinical experts (Appendix B) checked each case to ensure anonymity, realism, and pedagogical usefulness. As shown in Figures 3 and 2, SPBench covers 14 medical fields.
+Quality Control To ensure accuracy and reliability, we scanned the books using Optical Character Recognition (OCR).
+Then, three senior medical students manually verified the extracted text and corrected typographical and punctuation errors.
+This step ensures the data faithfully matches the original books, which is necessary for a reliable benchmark.
 Figure 3: Distribution of clinical cases in the SPBench dataset by medical department.
-Case Source and Scope
-We collected 58 real clinical cases.
-From the doctor-patient dialogue records associated with these cases, we processed and extracted a dataset containing 3,208 question-and-answer pairs.
-The source materials include the Manual for Writing Standardized Patient Cases [^1] https://weread.qq.com/web/bookDetail/1a432e10811e38145g0194ab and the A Practical Tutorial for Standardized Patients [^2] https://www.pumpedu.com/home-shop/7125.html.
-All records were fully anodized and used solely for academic research.
-They were also reviewed by experienced clinicians (see Appendix 9 for expert qualifications) to ensure accuracy and representativeness.
-As shown in Figure 2 and Figure 3, the specialty distribution of SPBench spans 14 clinical departments.
-General Surgery is the most represented specialty, accounting for 13.79% of the cases, followed by other departments such as Cardiology, Respiratory Medicine, and Psychiatry.
-The age of the patients ranges from children to the elderly, with the majority concentrated in the 21-60 age group.
-Regarding gender, the dataset is composed of 35 female cases, representing 60.3% of the total, and 23 male cases, representing the remaining 39.7%.
-Figure 3: Demographic distribution of cases in the SPBench dataset. The left panel shows the age distribution, and the right panel shows the gender distribution.
-Data Structure
-SPBench provides two core types of content: 1) Case Scripts: These provide standardized patient histories and physical examination findings for each case; and 2) Authentic Dialogue Records: These are turn-by-turn transcripts of conversations between physicians and human SP.
-The records serve as a benchmark to compare the conversational quality of an LLM-based virtual SP against a human SP for the same case.
-Quality Control
-To ensure accuracy and reliability, all dialogue transcripts were digitized from the two books in Section 3.2 using Optical Character Recognition (OCR) and then manually proofread by three fifth-year medical students.
-They corrected OCR-induced errors such as typos and punctuation to ensure fidelity to the original source material.
-This process preserves the authenticity and consistency required for benchmarking.
 4.3 Evaluation Protocol
-To systematically assess the performance of LLM as Standardized Patients, we (i) define what to measure via eight core evaluation criteria; (ii) validate the reliability of our automated scoring against clinical experts; (iii) conduct fair comparisons between human SP, EasyMED, and an external framework under identical inputs and scoring; and (iv) perform component-level ablations within EasyMED to analyze where gains come from.
-The following subsections detail each component.
-Evaluation Metric
-In collaboration with three clinical experts (see Appendix 9 for qualifications), we defined eight core dimensions for evaluating the virtual patient’s response quality, as detailed in Table 3.
-During the evaluation, experts rated performance for each dimension on a 5-point scale.
-These scores were then linearly transformed to a 100-point scale for reporting to enhance the granularity and readability of the results.
-Expert Rating and Correlation Analysis
-Considering the high cost of large-scale manual annotation, we first used a well-designed evaluation prompt (Appendix I.4) and employed GPT-4o as an automated evaluator to score all results.
-To verify its reliability (see Appendix 9 for qualifications), we randomly selected 30 samples for blind review by two clinical experts who were unaware of the GPT-4o ratings.
-The automated scores correlated strongly and significantly with the experts’ average ratings (Pearson’s $r=0.79$; Table 2).
-Although GPT-4o mean score (78.5) was consistently stricter than the experts’ mean (83.9), this high alignment supports its use as a reliable proxy for large-scale evaluation.
-Automated Professional A Professional B
-Average Score 78.5 86.3 81.5
-Standard Deviation 8.8 9.3 9.8
-Correlation 0.79
-Table 2: Inter-Rater Reliability between an Automated Evaluator and Human Experts
-Evaluation Criterion Description
-Query Comprehension Assesses the ability to accurately understand the physician’s questions, grasp the core intent, and provide relevant answers without misinterpretation.
-Case Consistency Evaluates whether responses are factually consistent with the predefined patient case, including symptoms, medical history, and key facts.
-Checks for contradictions or information not present in the case.
-Controlled Disclosure Measures the ability to provide information passively and appropriately, answering only what is asked without volunteering unsolicited details.
-Focuses on conciseness and relevance.
-Response Completeness Assesses whether responses are comprehensive and address all parts of a physician’s query, avoiding the omission of crucial information from the case profile.
-Logical Coherence Evaluates the internal logical consistency of the information provided.
-Ensures that reported symptoms and their characteristics (e.g., location, duration) are coherent from a patient’s perspective.
-Language Naturalness Assesses whether the communication style is natural and patient-like, using plain language and avoiding overly technical medical jargon.
-Focuses on the authenticity of the linguistic style.
-Conversational Consistency Measures the ability to maintain consistent information across a multi-turn dialogue. Checks for self-contradictions and ensures stability throughout the interaction.
-Patient Demeanor Evaluates the ability to maintain a patient-like demeanor, demonstrating appropriate patience, cooperation, and emotional tone, even when faced with repetitive or clarifying questions.
-Table 3: Definitions of the eight evaluation criteria for assessing the quality of the conversational agent.
-Comparison between human and LLM-based SP
-To compare the responses of LLM-based SP with those of a human SP under identical context, we first isolated the physician’s questions from the dialogue records in SPBench.
-We then used these exact questions and patient case as inputs for the LLM-based SP.
-This method allowed us to directly compare the responses generated by our virtual patients against those of the human SP under identical conditions.
-Comparison with External Frameworks
-Beyond comparing against human SP under identical prompts, we further benchmarked our system against an open-source LLM-based SP framework, EvoPatient du2024llms.
-Using the same SPBench physician questions and case profiles, and the same automated evaluator and scoring protocol, basic settings are summarized in Appendix F.
-we report the system-level scores in Table 4.
-System Configuration Score
-Human SP 97.33
-EasyMED (ours) More Agent 96.98
-EvoPatient Authors setup 93.33
-Table 4: Comparison of different frameworks on SPBench under the same evaluation protocol.
-Ablation on EasyMED Components
-We conducted an ablation study to quantify the contribution of key components in our virtual SP design.
-For each model, we started from a baseline configuration and incrementally added components: (1) Baseline no chain-of-thought or intent recognition; (2) CoT adds explicit step-by-step reasoning to the baseline; (3) Intent Recognition further adds an intent-recognition module on top of CoT.
-Model Name Baseline COT Intent Recognition
-Qwen3-8B 81.12 81.95 81.64
-Qwen3-32B 88.87 91.89 94.08
-DeepSeek-R1 93.11 94.90 95.72
-GPT-4o 89.54 90.08 93.70
-Gemini 2.5 Pro 95.04 95.77 96.98
-Table 5: Effect of CoT and intent recognition modules across different models.
-Table 4 shows that EasyMED matches human SP performance (96.98 vs.
-97.33) and clearly outperforms the external framework EvoPatient (93.33).
-These results indicate that our multi-agent design and evaluation-aligned prompting produce higher-fidelity SP behaviors under identical inputs and scoring.
-As shown in Table 5, adding COT to the baseline improves most models, and layering intent recognition on top yields further gains for medium-to-large models (e.g., Qwen3-32B, DeepSeek-R1, GPT-4o, Gemini 2.5 Pro).
-However, these advanced components are capacity-sensitive: for a small model such as Qwen3-8B, intent-guided CoT can introduce over-constraints and slightly reduce the score relative to plain CoT.
+Evaluation Metric SPBench evaluates VSP by assessing both turn-level response quality and session-level interactional behavior.
+Unlike existing benchmarks that rely on static knowledge tests or single aggregate scores, our evaluation decomposes patient performance into multiple clinically interpretable dimensions.
+Specifically, we define eight evaluation criteria (Table 2) in collaboration with three clinical experts (see Appendix B).
+These criteria capture complementary aspects of patient simulation, including accurate understanding of clinical inquiries, adherence to the predefined case, controlled disclosure of information, and consistency across dialogue turns.
+Each criterion is independently rated on a 5-point Likert scale, and scores are linearly rescaled to a 100-point scale for reporting.
+Input Standardization To ensure fair and reproducible evaluation, SPBench uses authentic questions extracted from real-world doctor–patient dialogues.
+For each clinical case, we isolate the sequence of questions asked by the human physician and present this exact sequence to each model.
+This design eliminates variability introduced by different prompting styles and ensures that all virtual standardized patients are evaluated under identical input conditions.
+5 Evaluation on SPBench
+This section analyzes divergences between virtual and human standardized patients in multi-turn interactions and assesses how EasyMED reduces these gaps relative to human SP behavior.
+Overall Performance  Table 3 summarizes overall and per-dimension performance on SPBench.
+Human standardized patients achieve the highest reference score (97.33), reflecting stable case portrayal and appropriate information disclosure across turns.
+EasyMED closely matches human performance (96.98), with strongest gains on interaction-critical dimensions (CC, CD, CS, PD) that directly align with standardized patient requirements.
+By contrast, although several frontier LLMs perform well on LN and RC, they show larger variance on CC and CD, indicating unstable case grounding and inconsistent inquiry-conditioned disclosure.
+Sensitivity on Prompting Strategies  To disentangle the sources of performance differences, we first examine LLM baselines under a unified prompting scheme.
+Although large models such as Gemini 2.5 Pro achieve relatively balanced scores, consistent weaknesses remain in controlled disclosure and cross-turn stability.
+We further evaluate common prompting strategies using a fixed backbone (Gemini 2.5 Pro).
+CoT prompting improves reasoning transparency and modestly increases CC and RC.
+However, it also amplifies verbosity and unsolicited explanation, leading to reduced CD scores.
+Overall, prompting mainly influences response articulation rather than information control, and is insufficient to enforce standardized patient behavior in multi-turn interactions.
+Ablation Study on Auxiliary Agent  We next examine the effect of the Auxiliary Agent in EasyMED, which decouples intent recognition from response generation.
+As shown in Table 3, this component yields the largest gains on CC, CD, CS, and PD.
+By mapping learner queries to standardized clinical intents, the Auxiliary Agent provides an explicit control signal that constrains information access, mirroring how human standardized patients condition responses on inquiry type rather than full case narratives.
+These gains exceed those from prompt engineering alone, highlighting the dominant role of architectural control in stabilizing multi-turn patient behavior.
+Figure 4: Learning trajectories of individual students and group averages for Group A (left) and Group B (right) across three assessment points.
+Each colored line tracks an individual student’s performance, while the bold dashed line represents the group’s average score.
+6 Real-world Evaluation in Medical Education
+While SPBench evaluates interaction-level fidelity, it does not directly capture educational effectiveness.
+We therefore assess EasyMED in a real training setting through a controlled user study, comparing it with human standardized patient training in terms of learning outcomes, learner experience, and practical feasibility.
+Study Period Timeline Group A Group B
+Baseline Week 0 Pre-Test
+Period 1 Weeks 1-2 EasyMED Training Human SP Training
+End of Week 2 Mid-Test
+Period 2 Weeks 3-4 Human SP Training EasyMED Training
+End of Week 4 Final Test & Questionnaire
+Table 4: Experimental design of the four-week study, showing the sequence of training interventions and assessments for Group A and Group B.
+6.1 Experimental Design
+We adopted a randomized crossover design to enable within-subject comparison while controlling for baseline differences.
+Each participant experienced both EasyMED and human SP training in different phases, which allows analysis of overall learning gains as well as phase-specific effects attributable to each modality (Table 4).
 Group Sequence Mean Score (±SD) at Test Point Mean Score Gain (±SD) by Phase
@@
 All values are mean $\pm$ standard deviation.
-4 Experiment Protocol
-After validating our virtual SP on SPBench, we conducted a user study to evaluate its educational impact in a real-world setting.
-This section describes the four-week randomized crossover trial with undergraduate medical students, designed to test whether EasyMED outperforms traditional human SP training.
-Study Period Timeline Group A Group B
-Baseline Week 0 Pre-Test Pre-Test
-Period 1 Weeks 1-2 Intervention A: Intervention B:
-EasyMED Training Human SP Training
-End of Week 2 Mid-Test Mid-Test
-Period 2 Weeks 3-4 Intervention B: Intervention A:
-Human SP Training EasyMED Training
-End of Week 4 Final Test & Questionnaire
-Table 7: Design of the Experiment, illustrating the sequence of interventions (EasyMED Training and Human SP Training) and assessments for Group A and Group B.
-4.1 Experimental Design
-We ran a four-week randomized crossover experiment comparing EasyMED with traditional human SP training.
-Each participant experienced both training methods in different phases, enabling within-subject comparisons and controlling for individual differences in baseline competency.
-The procedure is shown in Table 7.
-Patient Cases
-During the intervention, students practiced with 20 standardized clinical cases covering core diseases at the scope and difficulty of the National Medical Licensing Examination (see Appendix D).
-4.2 Participants
-We recruited 20 fourth- and fifth-year undergraduate medical students, all of whom completed a pre-test to assess baseline knowledge and received a 10-minute EasyMED orientation.
-Students with scheduling conflicts or extreme pre-test scores were excluded, resulting in a final sample of 14 (7 males, 7 females; mean age 23, range 21–-24).
-All had completed core clinical coursework but had not yet taken the National Medical Licensing Examination.
-To ensure baseline equivalence, participants were rank-ordered by pre-test scores and alternately assigned to two sequence groups.
-Three experienced SP professionals delivered the human SP sessions.
-All participants received hourly compensation in accordance with ethical regulations.
-5 Results and Analysis
-This section presents and analyzes the results from Section 4, focusing on whether the Virtual SP-based EasyMED can deliver learning outcomes equivalent to or better than traditional human SP training.
-5.1 Evaluating Overall Improvement via OSCE
-Figure 4: Boxplots comparing the baseline scores of Group A and Group B.
-Each circle represents the score of an individual participant.
-The similarity in the distributions demonstrates that the two groups were comparable at the start of the experiment.
-Figure 5: Learning trajectories of individual students and group averages for Group A (left) and Group B (right) across three assessment points.
-Each colored line tracks an individual student’s performance, while the bold dashed line represents the group’s average score.
-Both groups started from an equivalent baseline We first confirmed that the two groups were comparable at baseline.
-Baseline equivalence.
-Figure 4 shows that the two groups had similar initial score distributions.
-An independent-samples t-test confirmed no significant difference between Group A (mean = 70.56) and Group B (mean = 69.84), t(12) = 0.16, p = 0.88, indicating comparable baselines.
-Overall Performance Improvement:
-Both training methods produced clear, comparable gains.
-Specifically, As shown in Table 6, Group A’s average score rose from 70.56 to 87.44 (+16.89), while Group B’s increased from 69.84 to 85.20 (+15.36).
-Figure 5 illustrates this upward trend across sessions.
-Figure 6: Comparison of mean score gains in Phase 1 for the Human SP and EasyMED training methods.
-The participants are stratified into low- and high-performing groups based on their pre-test scores.
-Error bars indicate the standard error of the mean.
-Improvement at Each Stage
-Most gains occurred in the first two weeks, followed by a slowdown in Weeks 3-–4.
-Table 6 shows that Group A improved by 15.51 points and Group B by 12.17 points in the first two weeks.
-In Weeks 3–4 the rate of improvement slowed markedly, indicating diminishing returns and underscoring the strong initial impact of both training methods.
-Figure 7: This chart shows the total score gain for each group, broken down by training phase.
-The total height of each bar represents the average total gain for that group.
-The stacked segments show the portion of the gain that occurred during Phase 1 (Weeks 1-2) and Phase 2 (Weeks 3-4).
+Participants We recruited 20 medical undergraduate students in their fourth or fifth year.
+All participants completed a pre-test and received a 10-minute introduction to EasyMED.
+Students with scheduling conflicts or anomalous test scores were excluded (Appendix G), yielding a final cohort of 14 students (7 men, 7 women; age range 21–24, mean 23).
+All participants had completed core clinical coursework but had not yet taken the National Medical Licensing Examination.
+Participants were ranked by pre-test scores and assigned to groups using an alternating allocation scheme.
+Three experienced professionals served as human standardized patients.
+All participants were compensated on an hourly basis in accordance with institutional ethical guidelines.
+6.2 Results and Analysis
+6.2.1 Evaluating Overall Improvement via OSCE
+Objective Structured Clinical Examination (OSCE) is a standardized, station-based clinical skills assessment.
+We first confirm that the two groups were comparable prior to the intervention.
+As shown in Figure 5, baseline OSCE score distributions did not differ significantly between Group A (mean = 70.56) and Group B (mean = 69.84; $t(12)=0.16$, $p=0.88$), indicating similar starting levels.
+Figure 5: Boxplots of baseline OSCE scores for Group A and Group B prior to the intervention. Each point represents an individual participant.
+Overall Performance Across the four-week study, both groups demonstrate substantial and comparable improvements in OSCE scores.
+As summarized in Table 5, Group A improved by 16.89 points on average, while Group B improved by 15.36 points.
+Figure 4 shows consistent upward trends across individual learners in both groups.
+Finding 1: These results indicate that EasyMED supports clinical skill acquisition at a level comparable to human standardized patient training.
 Figure 6: Comparative heatmap of the weekly practice time distribution for the EasyMED and Human SP groups.
@@
 The color intensity and the white number in each cell indicate the number of students who practiced during that time slot.
-Improvement by Skill Level
-EasyMED delivered the largest benefits for low-baseline students early on.
-We split participants into high-baseline (top three) and low-baseline (bottom four) subgroups based on pre-test scores.
-As shown in Figure 6, low-baseline students using EasyMED gained an average of 21.83 points versus 16.58 in the human SP group.
-High-baseline students improved more modestly—7.10 for EasyMED versus 6.30 for human SP—showing that EasyMED is particularly advantageous for learners starting at a lower skill level.
-Improvement by Training Method
-Across the crossover design, EasyMED produced larger score gains than traditional SP training.
-Although the final post-test scores for Group A (M = 87.44) and Group B (M = 85.20) were not significantly different ($p>.05$), the crossover design allowed us to analyze the score gains from each training method independently (Figure 7).
-Training with EasyMED led to an average gain of 9.35 points, compared with 6.77 points for training with a human SP.
-This indicates that, while both methods are effective, EasyMED yields greater skill improvement than traditional SP training.
-5.2 Behavioral Analysis via Survey and Logs
-To investigate why EasyMED is effective, we analyzed both behavioral logs and questionnaire responses.
-Training Flexibility
-Students rated EasyMED more usable and used it far more flexibly, with practice extending into evenings and weekends.
-Students gave EasyMED significantly higher usability scores than human SP (4.8 vs.
-3.5, $p<.01$).
-This preference was reflected in activity patterns (Figure 8): human SP sessions clustered during weekday working hours, whereas EasyMED practice was spread throughout the day and peaked on evenings and weekends.
-These patterns indicate that EasyMED overcomes the scheduling limits of traditional training, offering more flexible and self-paced learning.
-Psychological Safety and Engagement
-EasyMED provided a lower-anxiety environment that encouraged deeper engagement and more questioning.
-Students reported significantly less learning anxiety with EasyMED than with human SP (mean score 2.5 vs.
+Phase-wise Effects Most learning gains occurred during the initial training phase for both modalities.
+During Phase 1 (Weeks 1–2), Group A gained 15.51 points using EasyMED, while Group B gained 12.17 points using human SPs.
+In Phase 2 (Weeks 3–4), when groups switched modalities, additional gains were observed but at a slower rate, suggesting diminishing returns commonly seen in short-term intensive training.
+Improvement by Skill Level To examine individual differences, we stratify participants into high-baseline (top three) and low-baseline (bottom four) groups based on their pre-test OSCE scores.
+As shown in Figure 7, low-baseline using EasyMED gained an average of 21.83 points, compared to 16.58 points with human SP.
+High-baseline improved less (7.10 vs.
+6.30).
+Finding 2: This indicates that EasyMED is particularly effective for novice learners during the early stage of training (i.e., the first two weeks).
+Figure 7: Comparison of mean score gains in Phase 1 for the Human SP and EasyMED training methods.
+The participants are stratified into low- and high-performing groups based on their pre-test scores.
+Error bars indicate the standard error of the mean.
+6.2.2 Behavioral Analysis via Survey and Logs
+To further contextualize the learning outcomes reported in Section 6.2.1, we analyzed students’ subjective questionnaire responses (see Appendix F) together with interaction logs collected during training.
+Perceived Authenticity Students reported high perceived realism when interacting with EasyMED.
+On a five-point Likert scale, the simulated patient dialogue achieved a mean authenticity score of 4.6 (Table 6), indicating that the interaction was generally regarded as natural and clinically plausible.
+Ratings for learning helpfulness were comparable to those of human SP training, suggesting that EasyMED is perceived not merely as a convenient substitute, but as a viable modality for practicing history-taking and clinical reasoning.
+Peer Pressure Survey results indicate substantially lower learning anxiety during EasyMED sessions than during human SP interactions (mean anxiety score 0.5 vs.
 3.2, $p<.01$).
-Interviews echoed this, with participants noting they were “not afraid to make mistakes” and felt comfortable “asking follow-up questions repeatedly.” This sense of safety translated into higher engagement: sessions with EasyMED averaged more dialogue turns (54 vs.
-47) and longer durations (28:49 vs.
-15:17) than human SP sessions (Table 8).
-Although part of the extra time reflects typing, the higher number of dialogue turns more strongly indicates engagement, suggesting that the low-anxiety environment promotes deeper reasoning and inquiry.
+Students reported feeling less concerned about making mistakes and more willing to ask exploratory or repeated questions.
+This low-pressure environment may facilitate risk-free exploration, particularly for learners at an early stage of training.
+Behavioral Evidence System logs provide objective evidence that complements these subjective reports.
+As shown in Figure 6, EasyMED practice sessions were distributed across a wide range of times, including evenings and weekends, whereas human SP sessions were largely confined to weekday working hours.
+In addition, EasyMED sessions involved more interaction on average, with a higher number of dialogue turns (54 vs.
+47) and longer session durations (28:49 vs.
+15:17) than human SP sessions (Table  6).
+Although text-based interaction may partially account for longer durations, the increased number of turns suggests more iterative questioning and sustained engagement.
 Metric EasyMED Human SP
 Student Engagement
-Learning Anxiety Score [^1] Anxiety was rated on a scale where lower scores are better. The difference is statistically significant ($p<.01$). 2.5 3.2
+Authenticity 4.6
+Helpfulness 4.5 4.7
+Learning Anxiety Score [^3] Anxiety was rated on a scale where lower scores are better. The difference is statistically significant ($p<.01$). 0.5 3.2
 Average Dialogue Turns 54 47
@@
 Table 6: Comparison of student engagement and cost-effectiveness metrics between EasyMED and human SP.
-Cost-Effectiveness:
-EasyMED delivers comparable training at a fraction of the per-session cost of human SP, making large-scale deployment feasible.
-Beyond its pedagogical benefits, EasyMED offers a substantial advantage in scalability and cost.
-As shown in Table 8, the per-session cost for an EasyMED interaction is approximately \$0.73, compared with \$52.95 for a human SP.
-This dramatic reduction in operational expense highlights the potential for EasyMED to provide high-quality clinical training at scale at only a fraction of the cost of traditional methods.
+Cost-Effectiveness: For cost estimation, EasyMED session costs are computed from the total token usage of a complete training interaction, whereas human SP costs are estimated by converting standard hourly compensation into a per-session cost.
+This results in an approximately 73-fold cost reduction compared to traditional SP training.
+Finding 3: EasyMED provides a realistic, low-pressure, and accessible training environment that supports sustained and exploratory practice at a fraction of the cost of human SP training.
 7 Conclusion
-This study examines the use of large language models in SP training.
-We present EasyMED, a multi-agent framework that separates patient simulation, intent recognition, and evaluation into coordinated agents aligned with real clinical workflows.
-To enable consistent evaluation, we build the SPBench benchmark from SP–student dialogues across 14 specialties and expert-defined criteria.
-A four-week controlled study shows that EasyMED matches human SP training in learning outcomes and improves performance for novice learners through higher flexibility, safety, and lower cost.
-Our work demonstrates the feasibility of multi-agent, LLM-based SP systems for scalable and effective clinical training.
-Future work will test larger cohorts, integrate multimodal interaction, and study long-term learning effects.
-7 Limitations
+This study examines whether large language models can function as standardized patients for clinical skills training.
+We propose EasyMED, a multi-agent framework for stable, inquiry-conditioned patient simulation, and introduce SPBench, a human-grounded benchmark built from standardized patient–student dialogues.
+In a four-week controlled study, EasyMED achieves learning outcomes comparable to human SP training, with stronger early gains for novice learners, greater flexibility, and substantially lower cost.
+These results suggest that LLM-based multi-agent VSPs are a practical and scalable complement to traditional SP programs.
+Limitations
 Our study has several limitations.
@@
 Future work will include larger multi-site and longitudinal studies, integration of multimodal interaction channels, and refinement of evaluation metrics to better capture nuanced performance.
-8 Ethical Statement
+Ethical Statement
 The study was approved by the institute on August 30, 2025. All annotators were fairly compensated, adhering to the standard hourly wage practices of their respective states.
 Appendix A Related Work
-This section reviews research on virtual patients and intelligent tutoring, the use of large language models (LLM) in education, and automated assessment of complex, interactive skills.
+This section reviews research on virtual patients and intelligent tutoring, the use of large language models in education, and automated assessment of complex, interactive skills.
 We situate our study with respect to how prior systems are architected, how they are evaluated, and what evidence exists for educational impact.
 A.1 Virtual Patients and Intelligent Tutoring
-Virtual patients (VP) have long supported safe practice of diagnostic reasoning and communication in medical education cook2009virtual; berman2016role; kononowicz2015virtual.
-Early systems were primarily rule- or script-based, providing structured but inflexible interactions and limited behavioral realism huang2007virtual.
-Recent work explores LLM-driven VP to increase linguistic fluency and adaptability du2024llms; almansoori2025self; grevisse2024raspatient; steenstra2025scaffolding.
+Virtual patients (VP) have long supported safe practice of diagnostic reasoning and communication in medical education Cook and Triola (2009); Berman et al.
+(2016); Kononowicz et al.
+(2015).
+Early systems were primarily rule- or script-based, providing structured but inflexible interactions and limited behavioral realism Huang et al.
+(2007).
+Recent work explores LLM-driven VP to increase linguistic fluency and adaptability Du et al.
+(2024); Almansoori et al.
+(2025); Grévisse (2024); Steenstra et al.
+(2025).
 However, most LLM-based VP adopt a paradigm that collapses patient simulation, dialogue control, and assessment into one model, typically focusing on the consultation/history-taking phase while leaving evaluation ad hoc and feedback coarse.
@@
 A.2 Large Language Models in Education
-LLMs have been studied for tutoring, content generation, and role-playing across domains gan2023large; wang2024large; xu2024large; safranek2023role.
-In medical education, they have been used to generate patient histories and to support clinical decision making, and to simulate patient dialogues for practicing history-taking waisberg2024large; thirunavukarasu2023large; almansoori2025self; fan2023nphardeval.
-Persistent challenges include factual reliability, long-context consistency, and alignment with professional standards and safety constraints li2024enhancing; mahaut2024factual; gai2024achieving.
+LLMs have been studied for tutoring, content generation, and role-playing across domains Gan et al.
+(2023); Wang et al.
+(2024); Xu et al.
+(2024); Safranek et al.
+(2023).
+In medical education, they have been used to generate patient histories and to support clinical decision making, and to simulate patient dialogues for practicing history-taking Waisberg et al.
+(2024); Thirunavukarasu et al.
+(2023); Almansoori et al.
+(2025); Fan et al.
+(2023).
+Persistent challenges include factual reliability, long-context consistency, and alignment with professional standards and safety constraints Li et al.
+(2024a); Mahaut et al.
+(2024); Gai et al.
+(2024).
 Most studies report surface metrics or static outcomes (e.g., diagnosis/referral accuracy) rather than interaction competencies.
@@
 A.3 Automated Assessment of Complex Skills
-Automated assessment with LLMs has advanced in essay scoring and feedback zhangpersonaeval; lee2024analyzing; gusev2024pingpong and program analysis for coding education parvathy2025automated; chen2020analysis, but much of this work provides single-dimensional scores and limited interoperability with respect to process.
+Automated assessment with LLMs has advanced in essay scoring and feedback Zhang et al.
+; Lee et al.
+(2024); Gusev (2024) and program analysis for coding education Parvathy et al.
+(2025); Chen et al.
+(2020), but much of this work provides single-dimensional scores and limited interoperability with respect to process.
 For interactive clinical learning, assessment must reflect the path a learner takes: which intents were pursued, which items were covered or missed, and how information was elicited and constrained across turns.
@@
 Appendix B Data Annotation Statement
-The annotation and validation of all datasets in this study were conducted by qualified medical professionals.
-Three clinical experts reviewed all cases to ensure medical accuracy and representativeness.
-Two licensed physicians with outpatient referral experience independently annotated and verified the dialogue data according to predefined guidelines, with only consensus cases retained.
-In addition, three fifth-year medical students performed manual proofreading and transcription of dialogue records to correct OCR errors.
-For the creation of standardized patient scripts, three SP education specialists collaborated with three professional standardized patient actors to design and validate clinically authentic dialogues.
-All contributors participated voluntarily and provided informed consent under institutional ethical approval.
+To ensure medical accuracy, pedagogical validity, and ethical compliance, we assembled a multidisciplinary team composed of clinical experts, licensed physicians, medical students, and standardized patient (SP) professionals.
+The specific roles and contributions were distributed as follows:
+Clinical Expert Panel
+A panel of three clinical experts, consisting of two senior physicians with eight years of clinical experience and one attending physician with five years of experience, was responsible for the high-level design and validation of the study.
+Their duties included defining the eight expert evaluation criteria for SPBench, validating the intent recognition dataset, and overseeing the selection of clinical cases from authoritative training sources.
+Data Annotation and Blind Review
+To ensure objectivity and inter-rater reliability, specific annotation tasks were conducted by two independent licensed physicians (with three and five years of clinical experience, respectively) who were blinded to the model outputs and student groupings.
+Their specific tasks included:
+  • Case Quality Control: Checking every processed case to ensure anonymity, realism, and teaching utility.
+  • Benchmarking: Conducting a blind review of 86 randomly selected samples to validate the automated GPT-4o evaluation scores.
+  • OSCE Scoring: Independently scoring the pre-, mid-, and post-experiment OSCE tests for all student participants.
+  • Evaluation Agent Validation: Assessing the clinical appropriateness and validity of the guidance generated by the Evaluation Agent across 30 distinct practice sessions (as detailed in Appendix  D).
+Data Processing and Annotation Support
+Three senior medical students (5th-year undergraduates) were recruited for data preparation tasks.
+  • Digitization: They performed manual proofreading and correction of OCR-scanned text from the source books to fix typos and punctuation errors.
+  • Auxiliary Agent Validation: They participated in the construction of the Intent Recognition Test Dataset (Appendix  C).
+This involved reviewing and filtering the preliminary corpus of clinical questions generated by GPT-4o to remove ambiguous or unrealistic entries, ensuring the dataset’s quality for testing the Auxiliary Agent.
+Standardized Patient Script and Simulation
+The creation of high-fidelity scripts and human SP performance involved a collaborative team of three SP education instructors (specializing in 5th-year medical student training) and three professional SPs (with two years of acting experience).
+This team designed the patient history, symptoms, and emotional cues.
+The same three experienced professionals served as the human SPs during the four-week comparative user study.
+Ethical Compliance
+All contributors, including students, actors, and experts, participated with informed consent. They were compensated for their time adhering to standard hourly wage practices.
 Appendix C Construction of the Intent Recognition Test Dataset
@@
 Expert Validation and Curation
-The preliminary corpus was subsequently reviewed by a three-person expert panel, composed of professional medical personnel.
+The preliminary corpus was subsequently reviewed by a three medical student panel, composed of professional medical personnel.
 The panel’s task was to remove any questions that were ambiguous, clinically unrealistic, or had unclear intent attribution to ensure the high quality and validity of each entry in the final dataset.
@@
 Gemini2.5-flash 96.3 96.1 93.9 95.0
-Table 10: Performance comparison of different models on the balanced intent recognition task, reflecting realistic fluctuations. All scores are percentages (%).
-Appendix D Clinical Case Data Preparation
+Table 8: Performance comparison of different models on the intent recognition task. All scores are reported as percentages (%).
+Appendix D Validation of the Evaluation Agent Guidance
+To ensure that the post-session guidance provided by the Feedback Agent (e.g., highlighting missed or superfluous inquiries) is clinically sound and pedagogically appropriate, we conducted an independent validation study.
+Study Design and Methodology
+We sampled 60 complete practice sessions from the user study.
+For each session, the specific feedback generated by the agent was extracted and anonymized.
+Two independent clinical experts, blinded to the source of the generation, rated the appropriateness of each feedback item on a 5-point Likert scale (1=misleading, 5=highly appropriate).
+We defined two primary evaluation metrics: Accuracy, calculated as the percentage of feedback items receiving a score of $\geq 4$ from both experts; and Inter-rater Agreement, measured using Cohen’s $\kappa$.
+Results and Discussion
+The validation results are summarized in Table 9.
+Across the 60 evaluations, the Feedback Agent demonstrated high reliability, achieving an accuracy of 87% with substantial agreement between experts ($\kappa=0.76$).
+Qualitative error analysis revealed that most disagreements arose in borderline cases where the clinical necessity of a specific inquiry was debatable.
+These findings confirm that the agent provides generally reliable guidance.
+Future iterations could incorporate confidence scores to allow experts to flag ambiguous feedback for refinement.
+Table 9: Validation results of the Feedback Agent’s guidance across 60 sessions. Accuracy is defined as the percentage of items rated $\geq 4$ by both clinical experts.
+Metric Value Description
+Sample Size 60 Total number of sessions evaluated
+Accuracy 87% Proportion of feedback rated as appropriate
+Inter-rater Agreement 0.76 Cohen’s $\kappa$ indicating expert consensus
+Appendix E Clinical Case Data Preparation
 This section details the source and selection criteria for the 20 clinical cases used in our user study, as well as the data preparation process for both the human SP and the EasyMED system.
-D.1 Case Source and Selection
+E.1 Case Source and Selection
 A panel of medical experts selected 20 clinical cases from the authoritative "Peking Union Medical College Hospital Clinical Thinking Training Case Collection," ensuring they were aligned with the curriculum for fifth-year undergraduate medical students.
@@
 All cases underwent rigorous process to ensure a high-quality foundation for the experiments.
-D.2 Data Processing Workflow
+E.2 Data Processing Workflow
 To support the EasyMED multi-agent framework in achieving high-fidelity SP simulations, we designed a two-stage data processing workflow: 1) generating performable dialogue scripts for human SP; and 2) structuring the clinical cases to be compatible with LLM inputs, aiming to reduce the risks of hallucination and irrelevant responses, and to support the full-workflow clinical simulation and automated evaluation.
@@
 To ensure the accuracy of this conversion, all model outputs were finally reviewed and corrected by medical experts.
-Appendix E Outcome Measures
+Appendix F Outcome Measures
 We collected data through both quantitative and qualitative methods. Our primary and secondary outcome measures are detailed below.
-E.1 Primary Outcome Measure: OSCE Scores
+F.1 Primary Outcome Measure: OSCE Scores
 We administered OSCE tests to all students at three time points: pre-experiment, mid-experiment, and post-experiment.
@@
 Scoring was performed independently by two blinded examiners who were unaware of the students’ group assignments, ensuring objectivity.
-E.2 Secondary Outcome Measure: Subjective Questionnaire
+All participants provided written informed consent prior to participation.
+F.2 Secondary Outcome Measure: Subjective Questionnaire
 At the end of the experiment, we used a subjective questionnaire with 25 items across four dimensions (Usability, Authenticity, Learning Value, and Learning Anxiety) to collect students’ perceptions and experiences of the two training modalities.
-E.2.1 Part 1: Background Information
+F.2.1 Part 1: Background Information
   1. 1.
@@
     (5-point scale: 1–Not interested at all / 2–Slightly uninterested / 3–Neutral / 4–Somewhat hopeful / 5–Very excited)
-E.2.2 Evaluation of Learning and Training Models
+F.2.2 Evaluation of Learning and Training Models
   1. 5.
@@
     (5-point scale: 1–Very unrealistic / 2–Unrealistic / 3–Neutral / 4–Realistic / 5–Very realistic)
-E.2.3 Overall Assessment and Open-ended Feedback
+F.2.3 Overall Assessment and Open-ended Feedback
   1. 22.
@@
     What do you think is the biggest advantage of learning with a Human SP? (e.g., emotional connection, non-verbal cues, etc.)
-Appendix F Experimental Settings
-F.1 A. EasyMED (ours)
+Appendix G Participant Exclusion
+We initially recruited 20 medical students.
+Before random group assignment and prior to any training, six participants were excluded based on predefined criteria, resulting in a final sample of 14 students.
+Specifically, four students were excluded due to scheduling conflicts that prevented them from attending the required in-person human SP sessions, and two students were excluded due to extreme pre-test OSCE scores (95 and 96 out of 100), where ceiling effects would limit measurable learning gains.
+All exclusions occurred before group assignment and independently of the intervention, ensuring no differential attrition between conditions.
+Although the excluded students had a higher mean baseline score than the included cohort, this does not introduce selection bias because exclusions were applied prior to randomization.
+Appendix H Experimental Settings
+H.1 A. EasyMED (ours)
 Backbone per agent. Patient Agent: Gemini2.5-pro; Auxiliary (intent) Agent: Gemini2.5-flask; Evaluation Agent: Gemini2.5-pro.
@@
 Session policy. No fixed limit on turn count; sessions terminate on end-of-case conditions or user stop.
-F.2 B. EvoPatient
+H.2 B. EvoPatient
 Backbone. Gemini2.5-pro.
@@
 Protocol parity. Same case pool and physician question lists as EasyMED.
-Appendix G User Interface of the EasyMED
+Appendix I User Interface of the EasyMED
 This section provides screenshots of the EasyMED virtual patient system’s user interface.
@@
 The system provides itemized feedback (e.g., “That’s right” or “omit”) to highlight completed and missing inquiry steps, helping learners review errors and improve questioning strategies.
-Appendix H Validation of the Feedback Agent’s Guidance
-H.1 Objective
-The feedback agent in our training platform provides post-session guidance to users, such as highlighting missed or redundant questions.
-To ensure that this guidance is appropriate and clinically sound, we conducted an independent validation study.
-H.2 Study Design
-  • Sample selection: We collected 30 complete practice sessions from participants using the platform.
-  • Feedback extraction: For each session, we extracted the feedback agent’s generated guidance, including “missed items” and “superfluous items.”
-  • Expert review: Two clinical experts, blinded to the feedback agent, independently rated the appropriateness of each feedback item on a 5-point Likert scale (1=misleading, 5=highly appropriate).
-H.3 Evaluation Metrics
-We computed:
-  • Accuracy of feedback: Percentage of feedback items judged "appropriate" (score $\geq$ 4) by both experts.
-  • Inter-rater agreement: Cohen’s $\kappa$ between the two experts for appropriateness labels.
-  • Error categorization: Manual classification of disagreements (false positives, omissions, ambiguity).
-H.4 Results Summary
-Across 30 sessions, 87% of the feedback agent’s suggestions were rated appropriate (score $\geq$ 4 by both experts).
-Inter-rater agreement between the two experts was $\kappa=0.76$.
-Most disagreements arose in borderline cases where the clinical relevance of a missed question was debatable.
-H.5 Discussion and Recommendations
-These findings indicate that the feedback agent provides generally reliable guidance but may require refinement for ambiguous or edge cases.
-Future work could incorporate a confidence score or allow experts to flag uncertain feedback for model retraining.
-All annotated feedback and rating sheets are included in the supplementary materials for replication.
-Appendix I Core Prompts Used in the Study
+Appendix J Core Prompts Used in the Study
 This section details the core prompts used for patient simulation and automated evaluation in our study.
-I.1 Patient Agent Prompt
+J.1 Automated Evaluation with GPT-4o
+To enable scalable and reproducible evaluation on SPBench, we employ GPT-4o as an automated judge to assess the quality of virtual standardized patient responses.
+This section details the evaluation pipeline, including model inputs, scoring procedure, and score aggregation.
+Evaluation Input.
+For each test instance, GPT-4o is provided with three components: (1) a structured case describing the ground-truth patient profile, (2) the full doctor–patient dialogue transcript, and (3) a fixed evaluation prompt specifying eight expert-defined evaluation criteria.
+The same prompt and input format are used for all evaluated systems to ensure consistency.
+Scoring Procedure.
+GPT-4o evaluates the patient responses independently along eight dimensions (Query Comprehension, Case Consistency, Controlled Disclosure, Response Completeness, Logical Coherence, Language Naturalness, Conversational Consistency, and Patient Demeanor).
+Each dimension is rated on a 5-point Likert scale following explicit rubric definitions.
+The model is instructed to justify each score by citing specific dialogue turns as evidence and to return the results in a structured JSON format.
+Score Aggregation.
+For reporting, raw Likert scores are linearly rescaled to a 0–100 scale.
+Dimension-level scores are averaged across all test dialogues, and an overall score is computed as the mean of the eight dimension scores.
+No manual intervention or post-hoc adjustment is applied during this process.
+Automated Professional A Professional B
+Average Score 84.1 91.3 86.4
+Standard Deviation 8.5 9.1 9.3
+Correlation 0.81
+Table 10: Inter-rater reliability between the automated evaluator (GPT-4o) and human clinical experts, measured by Pearson correlation
+LLM vs. Human Evaluation
+Considering the high cost of large-scale manual annotation, we used GPT-4o as an automated judge.
+We gave it a carefully written prompt (Appendix J.5).
+To verify its reliability, we randomly selected 86 samples for blind review by two clinical experts (see Appendix  LABEL:sec:_partner for rater qualifications) who were unaware of the GPT-4o ratings.
+The automated scores correlated strongly and significantly with the experts’ average ratings (Pearson’s $r=0.81$; Table 10).
+Although GPT-4o’s mean score (84.1) was consistently lower than the experts’ mean scores (88.9), this high alignment indicates that GPT-4o provides a reliable proxy for large-scale evaluation.
+J.2 Patient Agent Prompt
 The Patient Agent is responsible for generating realistic patient responses in simulated medical consultations. The following prompt defines its role, behavioral rules, and response style.
@@
 Medical Case Information Conversation History
-I.2 Auxiliary Agent Prompt
+J.3 Auxiliary Agent Prompt
 The following prompt defines the behavior of the Auxiliary Agent, which performs intent recognition during doctor–patient dialogue.
@@
 (The latest utterance to be classified.)
-I.3 Evaluation Agent Prompt
+J.4 Evaluation Agent Prompt
 The following prompt defines the behavior of the Evaluation Agent (Clinical Skills Evaluator), which assesses students’ performance against expert standard answers.
@@
 Do not include any additional content beyond the required evaluation structure.
-I.4 Automated Evaluation Prompt
+J.5 Automated Evaluation Prompt
 This prompt instructs the Evaluation to act as a professional medical dialogue evaluator, scoring each conversation along eight dimensions and returning structured JSON outputs for interpretability.